### v1

# Electrocardiogram Instruction Tuning for Report Generation

Abstract
Electrocardiogram (ECG) serves as the primary non-invasive diagnostic tool for cardiac conditions monitoring, are crucial in assisting clinicians. Recent studies have concentrated on classifying cardiac conditions using ECG data but have overlooked ECG report generation, which is not only time-consuming but also requires clinical expertise. To automate ECG report generation and ensure its versatility, we propose the Multimodal ECG Instruction Tuning (MEIT) framework, the first attempt to tackle ECG report generation with LLMs and multimodal instructions. To facilitate future research, we establish a benchmark to evaluate MEIT with various LLMs backbones across two large-scale ECG datasets. Our approach uniquely aligns the representations of the ECG signal and the report, and we conduct extensive experiments to benchmark MEIT with nine open source LLMs, using more than 800,000 ECG reports. MEIT’s results underscore the superior performance of instruction-tuned LLMs, showcasing their proficiency in quality report generation, zero-shot capabilities, and resilience to signal perturbation. These findings emphasize the efficacy of our MEIT [^1] All data and code will be available upon acceptance. framework and its potential for real-world clinical application.
Machine Learning, ICML
[…]
${}^{\text{1}}$he Ohio State University ${}^{\text{2}}$niversity of North Carolina at Chapel Hill ${}^{\text{3}}$mperial College London
${}^{\text{4}}$ King’s College London  ${}^{\text{5}}$ ${}^{\text{6}}$
$*{}^{*}$orrespondence to: Zhongwei Wan<wan.512@osu.edu>Mi Zhang<mizhang.1@osu.edu>
${}^{\text{\textdagger}}$ore contribution  ${}^{\ddagger}$ore advising
1 Introduction
Diagnosing heart diseases is of paramount importance, requiring the analysis of diverse multimodal data sources, which includes electrocardiogram (ECG) recordings and clinical reports. Cardiologists interpret these ECG recordings to generate comprehensive ECG reports for heart disease diagnosis, which is a complex and time-consuming process. Recently, artificial intelligence (AI) models have been used to facilitate ECG data analysis by categorizing ECG data (Chen et al., 2022a; Hu et al., 2023; Liu et al., 2023a; Li et al., 2023b). Despite these early efforts, the automatic generation of reports from ECG recordings remains an underexplored area. Unlike other AI-empowered medical report generation applications (e.g., radiology report), the primary challenges for ECG report generation stem from the distinct nature of ECG content. ECG reports, often comprising brief phrases that summarize signal patterns, contrast with detailed anatomical descriptions in radiology reports. This difference in content and semantic interpretation between imaging and signal data complicates the direct application of radiology-focused methods to ECG reports. Furthermore, there is still a lack of comprehensive benchmarks for evaluating the performance of ECG report generation.
To tackle these challenges, we introduce MEIT, a Multimodal ECG Instruction Tuning framework that extends the capabilities of LLMs in the cardiology context to generate ECG reports using ECG signals based on human instructions. Inspired by the versatility of LLMs (Achiam et al., 2023; Touvron et al., 2023b; Wan et al., 2023; Wang et al., 2024) in handling diverse language tasks simultaneously, we develop a specialized instruction tuning process for ECG report generation. MEIT aligns human instructions with ECG signals, enabling LLMs to generate professional-grade reports and exhibit zero-shot report generation capabilities under domain transfer scenarios across various continents and data collection devices. Specifically, leveraging publicly available ECG datasets, we construct a multimodal instruction dataset including ECG records, human instructions, and paired reports. Then, we propose an effective and efficient attention-based fusion method to integrate ECG and text representations in the latent space, enabling LLMs to understand signals for report generation without introducing additional training parameters in the attention layer.
In addition to the ECG report generation approach within MEIT, we introduce a comprehensive benchmark for ECG report evaluation, utilizing two datasets with 20K and 800K ECG-report pairs, respectively, across three evaluation tasks: report generation quality, zero-shot learning across datasets, and robustness analysis in the face of ECG signal perturbation. Utilizing the ECG report evaluation benchmark, we assess the proposed approach across nine open-source LLMs. The results demonstrate (1) the superior performance of MEIT in ECG report generation and the effective learning and alignment of ECG representations; (2) the effective transferability of LLMs under the MEIT framework in domain transfer scenarios.
To summarize, our primary contribution is the MELT framework, a novel approach to automating ECG report generation and evaluation based on LLMs. This framework incorporates a lightweight, attention-based fusion module effective across various LLM models. Furthermore, we establish a new benchmark for ECG report generation, which contains three aspects of the evaluation tasks. Our evaluations showcase the enhanced capabilities of instruction-tuned LLMs in generating ECG reports, highlighting the transferability in zero-shot test and robustness against data perturbations. The MEIT framework paves the way for future advances in automated ECG report generation and methodological innovations in integrating biomedical signals into LLMs.
2 Preliminaries
Electrocardiogram (ECG) measures the electrical activity of an individual’s heart over time. An ECG recording typically contains a 12-lead multivariate time series, which act as a 12-dimensional sequence of embeddings. The ECG signal offers a comprehensive view, encompassing both spatial and temporal aspects of cardiac function. ECG leads can be categorized into 6 limb leads (i.e., I, II, III, aVR, aVL, and aVF) to monitor arms and legs, providing frontal plane views, and 6 precordial leads (i.e., V1, V2, V3, V4, V5, and V6) to monitor chest, showing horizontal plane views. We denote an ECG recording as $\mathbf{X}_{e}\in\mathcal{R}^{M\times T}$, where $M$ represents the numbers of leads, $T$ is signal length. Each ECG recording is associated with an ECG report $\mathbf{X}_{t}$ for description. Thus, we denote each ECG pair as $\{\mathbf{X}_{e},\mathbf{X}_{t}\}$. More visualization details can be found in the Appendix.
Figure 1: Illustration of the proposed MEIT framework, including ECG instruction data curation, instruction tuning on auto-regressive objective, and inference to get the generated report.
Figure 2: Illustration of model architecture for ECG report generation. $\mathbf{K}_{e}$ and $\mathbf{V}_{e}$ refer to linear projection of $\mathbf{H}_{e}$ by multiplying shared $\mathbf{W}_{k}$ and $\mathbf{W}_{v}$ in the attention layer.
3 MEIT
Figure 1 illustrates the proposed MEIT framework. We begin by describing the curation process for ECG instruction data. We then delve into the architecture of the multimodal LLM designed for ECG report generation. Lastly, we describe the approach used to generate ECG reports with instruction tuning.
3.1 ECG Instruction Data Curation
Given an ECG signal $\mathbf{X}_{e}$, our goal during inference is to generate an ECG report using an instruction prompt. For instance, the prompt might be, “Given the ECG signal embeddings, please help me generate an accurate description for this ECG signal embeddings: ”. To achieve this goal, during the training phase, we aim to create instruction tuning data to generate a response $\mathbf{\hat{X}}_{t}$ that aligns semantically with the ground truth $\mathbf{X}_{t}$. In addition, since we cannot predict the exact instruction prompt that users will use, we need to ensure that our report generation process is robust enough to handle different prompts. To address this challenge, we first manually design some prompt samples, then utilize GPT-4 (Achiam et al., 2023) to generate a set of prompts by rephrasing, as shown in Figure 1. Then, we randomly select one instruction prompt $\mathbf{X}_{p}$ from the prompt set, and create a general instruction-following template: <|user|>: $\mathbf{X}_{p}$, $\mathbf{X}_{e}$}<|assistant|>: $\mathbf{X}_{t}$} </s>, where <|user|> and <|assistant|> are added special tokens for tokenizer, </s> is a stop sign for each response. This approach ensures that the generated response not only conveys the same meaning as the ground truth but also remains adaptable to different instruction prompts. Following this strategy, we construct the ECG instruction data using a MIMIC-IV-ECG (Gow et al., ) dataset that contains 80K annotated data, and a 20K dataset PTB-XL (Wagner et al., 2020).
3.2 ECG Report Generation Model
In MEIT framework, the multimodal ECG report generation model aims to decode ECG signals end-to-end and accurately generate ECG reports. The architecture is illustrated in Figure 2. Specifically, the report generation model aims to encode the ECG-signal $\mathbf{X}_{e}\in\mathcal{R}^{M\times T}$ into latent embeddings and integrate them with the language embeddings with modality alignment, and then autoregressively generate the ECG report. In the following, we detail each component of the report generation model.
ECG Encoder. Since ECG signal is of high resolution in the temporal domain, it is vital to efficiently extract temporal features per lead before interaction with semantic embeddings inside the LLM backbone. Our default ECG encoder $\mathcal{F}_{e}(\cdot)$ consists of some temporal convolution blocks to encode each ECG signal into embeddings. Specifically, each temporal convolution block is composed of several 1-D convolution layers, batch normalization layers, and ReLU activation layers followed by average pooling. To align the output dimension with the head dimension of LLM backbone $\mathcal{F}_{l}(\cdot)$, we employ a non-linear up-to-down projection layer $\mathcal{P}_{e}(\cdot)$ to obtain the ECG embeddings:
$\mathbf{H}_{e}=\mathcal{P}_{e}\left(\mathcal{F}_{e}\left(\mathbf{X}_{e}\right)\right),$ (1)
where $\mathbf{H}_{e}\in\mathcal{R}^{D_{h}}$, $D_{h}$ has the same dimension as the multi-head attention layers of LLMs. Note that our default ECG encoder is lightweight and is able to learn temporal patterns of signals without a long training period.
ECG Modality Alignment. To guide the LLMs in aligning ECG signal data with corresponding textual outputs, we introduce an ECG modality alignment strategy. This approach is detailed in Figure 2. Specifically, given the ECG embeddings $\mathbf{Z}_{e}$, the alignment strategy aims to incorporate $\mathbf{Z}_{e}$ with the current hidden state $\mathbf{H}_{t}^{i}$ generated from previous $i-1$-th layer of the LLM backbone $\mathcal{F}_{l}\left(\cdot\right)$ for next-token prediction task. Here $\mathbf{H}_{t}^{i}$ is defined as:
$\mathbf{H}_{t}^{i}=\mathcal{F}_{l}^{i-1}\left(\left[\mathbf{X}_{p},\mathbf{X}_{t}\right]\right),$ (2)
where $i$ is the current layer index. Different from gated-attention fusion in Flamingo (Alayrac et al., 2022) and Memorizing Transformer (Wu et al., 2022), or Q-former in BLIP-2 (Li et al., 2023a) that require the additional trainable parameters or design complex modules, we propose a simple yet effective concatenated-fusion method that directly injects the ECG embeddings with language context in the causal-attention (C-Attn) of the models. In our approach, each attention layer combines $\mathbf{H}_{e}$, generated from the ECG encoder and projector as a prefix condition, with $\mathbf{H}_{t}^{i}$, derived from the preceding layer. The fusion process is as follows:
$\operatorname{C-Attn}\left(\mathbf{H}_{e},\mathbf{H}_{t}^{i}\right)=\left[\operatorname{head}_{1},\ldots,\operatorname{head}_{k}\right]\mathbf{W}_{o},$ (3)
here, $k$ represents the number of attention heads, and $\mathbf{W}_{o}$, a matrix in $\mathcal{R}^{kD_{h}\times D_{m}}$, serves as the projection matrix with $D_{m}$ denoting the hidden size of the LLM backbone. We replicate $\mathbf{H}_{e}$ for each head $k$ times, merging the ECG and language features in the sequence dimension. This is achieved through a shared projection of keys and values for each pattern. The fusion is then articulated as:
$\mathbf{K}_{m,j}=[\mathbf{K}_{e,j},\mathbf{K}_{t,j}]^{\top},\mathbf{V}_{m,j}=[\mathbf{V}_{e,j},\mathbf{V}_{t,j}],$ (4)
$\operatorname{head}_{j}=\operatorname{Softmax}\left(\frac{\mathbf{Q}_{t,j}\mathbf{K}_{m,j}}{\sqrt{D_{h}}}\right)\mathbf{V}_{m,j},$ (5)
where $\mathbf{Q}_{t,j}=\mathbf{H}_{t,j}^{i}\mathbf{W}_{q,j}$, $\mathbf{K}_{e,j}=\mathbf{H}_{e}\mathbf{W}_{k,j}$, and $\mathbf{K}_{t,j}=\mathbf{H}_{t,j}^{i}\mathbf{W}_{k,j}$, with a similar notation for $\mathbf{V}_{e,j}=\mathbf{H}_{e}\mathbf{W}_{v,j}$ and $\mathbf{V}_{t,j}=\mathbf{H}_{t}^{i}\mathbf{W}_{v,j}$. Concatenation is denoted by []delimited-[][$\cdot$][ ⋅ ], and $\mathbf{K}_{m,j}$ and $\mathbf{V}_{m,j}$ symbolize the amalgamated features of query and key. $\mathbf{W}_{q,j}$, $\mathbf{W}_{k,j}$, and $\mathbf{W}_{v,j}$ in $\mathcal{R}^{D_{h}\times D_{h}}$ represent the projection matrices for query, key, and value for each head $j$, respectively. Our model’s design allows for the efficient fusion of two modalities through causal attention, facilitating conditional generation without the need for additional parameter updates to align the ECG modality.
3.3 ECG Instruction Tuning
As described in Section 3.1, we have converted ECG-text pairs into a chat-bot style instruction format: <|user|>: $\mathbf{X}_{p}$, $\mathbf{X}_{e}$}<|assistant|>: $\mathbf{X}_{t}$}</s>. During instruction tuning, we compute autoregressive loss only on tokens after response tokens <assistant>, and use label loss masking to finetune the model, where we mask all tokens belonging to $\mathbf{X}_{p}$ and $\mathbf{X}_{e}$. To save computational resources and accelerate the convergence of instruction tuning, we use the LoRA (Hu et al., 2021) adapters for all linear layers of the LLM backbone $\mathcal{F}_{l}$ and freeze its backbone. Subsequently, given a sequence of ECG instruction data, we compute the probability of the target response $\mathbf{X_{t}}$ as an autoregressive function:
$p\left(\mathbf{X}_{t}\mid\mathbf{X}_{p},\mathbf{X}_{e}\right)=\prod_{i=j}^{L}p_{\boldsymbol{\theta}}\left(\mathbf{x}_{t,i}\mid\mathbf{X}_{p},\mathbf{X}_{e},\mathbf{X}_{t,<i}\right),$ (6)
where $j$ is the start index after <assistant>, $\theta$ is the trainable parameters of LoRA and ECG encoder $\mathcal{F}_{e}$, $\mathbf{X}_{t,<i}$ is the response tokens before the current generation $\mathbf{x}_{t,i}$.
4 Benchmarking ECG Report Generation
Table 1: Natural language generation metric on MIMIC-IV-ECG. For model size, ’M’ denotes the million level, and ’B’ denotes the billion level. † refers to an improved instruct fine-tuned version of models. The light teal color indicates the second highest results, and heavy teal color indicates the highest results.
Models Size BLEU-1 BLEU-2 BLEU-3 BLEU-4 METEOR ROUGE-L ROUGE-1 ROUGE-2 CIDEr-D
GPT2-Medium 345M 0.576 0.527 0.456 0.425 0.551 0.523 0.544 0.512 3.70
GPT2-Large 774M 0.614 0.563 0.490 0.476 0.595 0.571 0.585 0.538 4.21
GPT-Neo 2.7B 0.631 0.579 0.534 0.489 0.727 0.689 0.715 0.592 4.81
GPT-NeoX 20B 0.645 0.588 0.539 0.523 0.719 0.701 0.712 0.622 4.92
GPT-J 6B 0.676 0.628 0.584 0.542 0.756 0.721 0.744 0.632 5.23
BLOOM 7B 0.669 0.624 0.591 0.550 0.758 0.725 0.745 0.639 5.19
OPT 6.7B 0.673 0.616 0.598 0.532 0.755 0.732 0.743 0.631 5.32
LLaMA-1 7B 0.685 0.648 0.615 0.543 0.761 0.724 0.742 0.642 5.26
Mistral 7B 0.697 0.659 0.611 0.571 0.763 0.740 0.765 0.658 5.48
LLaMA-2† 7B 0.706 0.662 0.622 0.581 0.775 0.745 0.768 0.664 5.55
Mistral-Instruct† 7B 0.714 0.665 0.619 0.576 0.768 0.751 0.762 0.667 5.62
Table 2: Natural language generation metric on PTB-XL. The light teal color indicates the second highest results, and heavy teal color indicates the highest results.
Models Size BLEU-1 BLEU-2 BLEU-3 BLEU-4 METEOR ROUGE-L ROUGE-1 ROUGE-2 CIDEr-D
GPT2-Medium 345M 0.329 0.278 0.254 0.232 0.441 0.391 0.561 0.433 2.12
GPT2-Large 774M 0.437 0.395 0.355 0.320 0.575 0.481 0.652 0.527 3.25
GPT-Neo 2.7B 0.474 0.449 0.398 0.373 0.602 0.486 0.674 0.595 3.70
GPT-NeoX 20B 0.469 0.453 0.417 0.399 0.620 0.553 0.688 0.622 3.58
GPT-J 6B 0.485 0.452 0.428 0.405 0.656 0.550 0.662 0.613 3.72
BLOOM 7B 0.491 0.462 0.427 0.415 0.665 0.580 0.678 0.605 3.80
OPT 6.7B 0.502 0.477 0.431 0.418 0.662 0.568 0.669 0.624 3.94
LLaMA-1 7B 0.514 0.485 0.465 0.430 0.678 0.588 0.682 0.613 3.97
Mistral 7B 0.486 0.475 0.446 0.421 0.673 0.591 0.697 0.634 3.98
LLaMA-2† 7B 0.515 0.484 0.469 0.439 0.675 0.594 0.698 0.624 4.05
Mistral-Instruct† 7B 0.501 0.481 0.457 0.425 0.664 0.592 0.700 0.641 4.01
Table 3: Semantic similarity between the generated ECG reports and ground truths is measured using BERTScore, denoted as P for Precision, R for Recall, and F-1 for the F-1 Score.
MIMIC-IV-ECG PTB-XL
Models P R F-1 P R F-1
GPT2-Medium 0.562 0.453 0.502 0.534 0.465 0.497
GPT2-Large 0.657 0.574 0.613 0.625 0.553 0.586
GPT-Neo 0.723 0.633 0.675 0.675 0.588 0.628
GPT-NeoX 0.719 0.638 0.676 0.654 0.579 0.614
GPT-J 0.725 0.655 0.688 0.689 0.622 0.654
BLOOM 0.734 0.684 0.708 0.701 0.645 0.672
OPT 0.713 0.667 0.689 0.712 0.648 0.678
LLaMA-1 0.752 0.697 0.723 0.725 0.657 0.689
Mistral 0.761 0.732 0.746 0.711 0.664 0.687
LLaMA-2† 0.764 0.725 0.744 0.721 0.668 0.693
Mistral-Instruct† 0.773 0.722 0.747 0.730 0.661 0.694
In this section, we create a benchmark to evaluate the ECG report generation approach. Our benchmark includes two datasets (PTB-XL and MIMIC-IV-ECG) along with three ECG-related tasks, including (1) the quality of ECG report generation; (2) the zero-shot ability under domain transfer; (3) the robustness of model against perturbed signals.
4.1 Datasets
PTB-XL. The PTB-XL dataset (Wagner et al., 2020) comprises 21,837 clinical 12-lead ECGs, each lasting 10 seconds, from 18,885 patients. Each ECG is accompanied by a corresponding report. The data includes raw ECG signals and their reports, with all recordings sampled at 500Hz over 10 seconds. We divided this dataset into training, validation, and testing subsets in a 70%:10%:20% ratio, respectively. The human experts double-check all samples in the test set to ensure data quality. We reformulate the dataset into instruction data format as mentioned in Sec 2.
MIMIC-IV-ECG. The MIMIC-IV-ECG dataset (Gow et al., ) is currently the largest publicly accessible ECG dataset, containing 800,035 paired samples from 161,352 unique subjects. Similar to PTB-XL, each sample in this dataset includes a raw ECG signal and its corresponding report, with all recordings sampled at 500Hz for 10 seconds. The division of this dataset into training, validation, and testing subsets is 80%:10%:10% ratio. Likewise, we reconstruct this dataset into an ECG instruction data template.
4.2 Evaluation Metrics
We evaluate models using nine metrics: BLEU 1-4 (Papineni et al., 2002), METEOR (Banerjee & Lavie, 2005), ROUGE 1-2 and L (Lin, 2004), CIDEr-D (Vedantam et al., 2015), and BERTScore (Zhang et al., 2019). BLEU and METEOR assess the quality of machine translation, focusing on accuracy and fluency. ROUGE-L measures sentence fluency and structure, while ROUGE-1 and ROUGE-2 examine uni-gram and bi-gram overlaps. CIDEr-D evaluates the relevance and uniqueness of generated ECG reports against a candidate set, and BERTScore assesses semantic similarity to the ground truth, ensuring content accuracy.
4.3 Tasks
We evaluate the performance of ECG report generation on the following three tasks:
Quality Report Generation. The goal of this task is to evaluate the quality of report generation following ECG instruction tuning. We utilize 10% of the MIMIC-IV-ECG and PTB-XL datasets, designated for instruction tuning, as the test set. This evaluation focuses on the syntactic and semantic similarities between the generated ECG report and the ground truth, based on diverse instructions and input ECG signals. For task assessment, we employ BLEU-1 to 4, METEOR, ROUGE 1, 2, and L, CIDEr-D, and BERTScore.
Zero-shot Generalizability. To explore the generalizability of LLMs in domain transfer scenarios following ECG instruction tuning, we trained the models on 70% of the instruction data from MIMIC-IV-ECG. We then assessed the zero-shot capabilities of these models on the PTB-XL test set, which contains high-quality ground truth evaluated by clinical experts. For model evaluation, we used the metrics BLEU-3, BLEU-4, METEOR, and ROUGE-L and calculated their average.
Signal Perturbation Robustness. In real-world clinical settings, ECG signals often contain some degree of noise. To evaluate the robustness of MEIT against such noisy interference, we selected 10% of the ECG samples from the MIMIC-IV-ECG test dataset. We then added Gaussian noise to these samples during the models’ instruction-based inference process. For this evaluation, we used BLEU-4, METEOR, ROUGE-L, and CIDEr-D as metrics.
5 Experiments and Evaluation
In this section, we evaluate and benchmark nine open-source decoder-only LLMs using the constructed ECG report generation benchmark. Additionally, we offer a comprehensive analysis of scalability and instruction tuning and present case studies showcasing the generated reports.
5.1 Experimental Settings
Models. We use nine LLMs based on the peft [^2] https://github.com/huggingface/peft library, which directly supports the LoRA (Hu et al., 2021) module, to construct the multimodal ECG report generation model described in Section 3.2. These models include GPT-Neo (Black et al., 2021), GPT-NeoX (Black et al., 2022), GPT-J (Wang & Komatsuzaki, 2021), BLOOM (Workshop et al., 2022), OPT (Zhang et al., 2022), LLaMA-1 (Touvron et al., 2023a), LLaMA-2 (Touvron et al., 2023b), Mistral (Jiang et al., 2023), and Mistral-Instruct [^3] https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.1, along with two relatively small pre-trained language models (GPT2-Medium and GPT-Large (Radford et al., 2019)) as fundamental baselines.
Implementation Details. Our experiment is implemented using PyTorch 2.1, the latest transformers (Wolf et al., 2020), and the accelerate (Gugger et al., 2022) library on 4 $\times$ A100 GPUs. We utilize the official weights for all LLMs from Hugging Face (Wolf et al., 2019) for evaluation, with sizes ranging from 2.7 billion to 70 billion parameters. For models exceeding 10 billion parameters, we use DeepSpeed [^4] https://github.com/microsoft/DeepSpeed for training. All models are trained on both MIMIC-IV-ECG and PTB-XL instruction datasets for 5 epochs with a learning rate of 2e-5 and a batch size of 64. We adopt a linear optimizer scheduler with a warm-up ratio of 0.03. The best performance of each model is determined by training on the training set, selecting from the validation set, and evaluating on the test set. More details about instruction tuning can be found in the appendix.
5.2 Benchmark Task Evaluation
5.2.1 Quality Evaluation
Performance on MIMIC-IV-ECG. Table 1 and 3 present the results of various types of language encoders $\mathcal{F}_{l}(\cdot)$ on MIMIC-IV-ECG. The results show that all LLMs exhibit superior performance compared to smaller language models (SLMs), such as GPT2-Medium and GPT2-Large, across all evaluation metrics. Notably, from GPT-Neo to Mistral-Instruct, LLM-based backbones achieve a significant margin over SLMs in all metrics. For instance, compared to GPT2-Large, the METEOR score increases in the range of 0.132 to 0.18 from GPT-Neo to LLaMA-2, and Mistral-Instruct outperforms GPT2-Large with an improvement of 0.18 in the ROUGE-L score and 0.134 in the F-1 of BERTScore.
The observed performance underscores the adeptness of LLMs in generalizing from signal data, showcasing enhanced proficiency in aligning ECG signal representations with corresponding textual information. This highlights the significant potential of LLMs in the domain of medical signal-to-text generation. Particularly, LLaMA-2 and Mistral-Instruct surpass their counterparts in the majority of evaluative metrics, suggesting that models pre-tuned with general instructions are more adept at learning ECG-text alignment.
Figure 3: Zero-shot performance on PTB-XL dataset. “IT” denotes instruction tuning.
Figure 4: Signal perturbation robustness analysis on various LLMs
Performance on PTB-XL. As shown in Table 2, the models exhibit reduced performance on PTB-XL compared to MIMIC-IV-ECG, which is attributable to the smaller scale of the instruction data in PTB-XL. This underscores the importance of data scale in enhancing instruction-based ECG report generation. Moreover, similar to the MIMIC-IV-ECG results, all LLM-based models show significant improvement over SLMs. Specifically, LLaMA-2 surpasses GPT2-Large by 0.134 in the BLEU-3 metric, while LLaMA-1 achieves a 0.103 improvement in the METEOR score. The overall experimental results reveal that Mistral-Instruct and LLaMA-2 are consistently the top two performers across most metrics, because of their strong general instruction-following capabilities. Consequently, we select LLaMA-2 and Mistral-Instruct as the default architectures for our framework.
5.2.2 Zero-shot Evaluation in Domain Transfer.
In Figure 3, we present the evaluation of the zero-shot learning capabilities of various LLMs, which is trained on the MIMIC-IV-ECG dataset and then tested on PTB-XL (unseen dataset). The assessed models include BLOOM, OPT, LLaMA-1, and Mistral. Firstly, all selected LLMs undergo instruction tuning on the MIMIC-IV-ECG train set, followed by zero-shot testing on the PTB-XL test set verified by human experts, denoted as Zero-shot IT. We also measure the performance of each model in report generation without prior ECG-specific instruction tuning, denoted as Zero-shot W/O IT. PTB-XL IT represents training on the PTB-XL train set and then evaluated on the PTB-XL test set. Notably, although Zero-shot IT shows a slight degradation compared to PTB-XL IT, the results still indicate a variance in the models’ ability to generalize to an unseen dataset with instruction tuning (IT), compared to Zero-shot W/O IT. The involvement of ECG instruction tuning on MIMIC-IV-ECG enables the models to achieve superior zero-shot performance on the unseen PTB-XL dataset, indicating the necessity of instruction tuning in enhancing the models’ zero-shot ability on unseen datasets in ECG report generation.
Figure 5: Ablation Study of ECG Instruction Tuning on MIMIC-IV-ECG Dataset.
5.2.3 Robust Analysis with Perturbed ECG Signal.
In a noise stress evaluation (Wang et al., 2019), we added Gaussian noise to ECG signals at signal-to-noise ratios (SNRs) of 0.05, 0.1, 0.15, and 0.2 during testing to assess model robustness. Our experiments utilized four LLM architectures: BLOOM, OPT, LLaMA-1, and Mistral, each trained on clean ECG signals from the MIMIC-IV-ECG training set and tested on corresponding noise-added signals from its test set. The results, illustrated in Figure 4, show a performance decline in all LLMs as SNR increased, highlighting the significant interference of ECG noise. Furthermore, as shown in Table 1, Mistral also excelled in tests on noise-free datasets, suggesting a synergistic effect between clean and noisy test sets. The results demonstrate Mistral’s strong resistance to perturbations. Even with more severe noise, it maintained robustness in terms of ROUGE-L and METEOR metrics. Developing an even more robust framework is a goal for future research.
5.3 Analysis
5.3.1 Instruction Tuning Visualization.
Figure 6 compares the convergence curves of the instruction tuning loss and the METEOR score between GPT-Neo (2.7B), BLOOM (7B), OPT (6.7B), and LLaMA-2 (7B) on the MIMIC-IV-ECG train and validation datasets. We observe that larger models with more parameters can converge to a smaller loss and achieve higher performance on the METEOR score. Notably, an increase in model size correlates with higher performance and lower loss, suggesting that larger models have the potential for better performance.
Figure 6: Instruction tuning loss and METEOR visualization.
Figure 7: Model Scaling Performance on MIMIC-IV-ECG and PTB-XL.
5.3.2 Scalability Analysis.
To investigate whether ECG instruction tuning on larger-scale models yields better results, we validated LLaMA-2 models of 7B and 13B parameter sizes on both MIMIC-IV-ECG and PTB-XL datasets. As depicted in 7, an upward trend in all evaluation metrics is observed with a gradual increase in model size. However, it is noteworthy that the gains in performance associated with increasing model size are not particularly significant. For example, the F-1 score for the 70B model on the PTB-XL dataset exhibits a marginal increase of 0.02 over the 13B model. Similarly, on the MIMIC-IV-ECG dataset, the 70B model’s F-1 score is only 0.01 higher than that of the 13B model. Therefore, we conjecture that enhancing both data scale and model size concurrently is necessary to achieve superior performance (Wei et al., 2022).
5.3.3 Ablation Study on ECG Instruction Tuning.
To validate the effectiveness of adopting an instruction tuning paradigm for aligning ECG signals and report representations, we conducted the following ablation experiment. We selected BLOOM, OPT, LLaMA-1, and Mistral as LLM backbones and removed the instruction tuning strategy, allowing the language models to fine-tune directly in the setting of generating text from ECG signals only. As shown in Figure 5, aligning ECG and report representations using direct fine-tuning, as opposed to instruction tuning, results in a significant decline across all metrics. Particularly in the Mistral model, the instruction tuning paradigm notably outperforms direct fine-tuning. These results demonstrate that, compared to fine-tuning, instruction tuning more effectively enhances the generalization ability of LLMs on unseen tasks/data (Ouyang et al., 2022).
Figure 8: Examples of ECG reports generated by LLaMA-2 and Mistral-Instruct. We highlight the consistent information between the generated reports and the ground truths with blue color.
5.3.4 Qualitative Results.
In Figure 8, we randomly select two samples generated by MEIT using LLaMA-2 and Mistral-Instruct as the LLM backbones. The consistent key information, highlighted in blue, indicates that both models have successfully learned important patterns from the ECG signals. Overall, the models’ results align with the ground truth, accurately identifying cardiac abnormalities from the ECG signals. Furthermore, both models provide detailed explanations of abnormal ECG signal details, such as ‘ischemia’ from sample 1 and ‘right bundle branch block’ from sample 2. These generated reports demonstrate the efficacy of our method.
6 Related Work
Instruction Tuning. Instruction tuning (Zhang et al., 2023a; Wang et al., 2023) is designed to enhance the zero-shot capabilities of LLMs on new tasks using either machine-generated or human-annotated instructions. Notable examples such as InstructGPT (Ouyang et al., 2022), FLAN-PaLM (Chung et al., 2022), and Alpaca (Taori et al., 2023) utilize various methods, including manual collection and human feedback, to generate instruction template data for fine-tuning. In the realm of multimodal learning, models like LLaVA (Liu et al., 2023b), MiniGPT-4 (Zhu et al., 2023), and AnyMAL (Moon et al., 2023) demonstrate the utility of multimodal instruction data to enhance learning across different modalities. However, medical signals such as ECG and their corresponding text remain underexplored. Our work bridges this gap by developing a specialized ECG-related instruction tuning framework, thereby improving the generation of ECG reports.
Medical Report Generation. Medical Report Generation can be categorized into three main approaches: (1) Template Selection and Generation, where methods like HRGR (Li et al., 2018) focus on sentence creation or template retrieval, and CMAS (Jing et al., 2017) uses dual writers for differentiating regions; (2) Data Integration and Coherence, illustrated by PPKED’s (Liu et al., 2021b) knowledge merging for consistency, and CA’s contrasting of abnormal features (Ma et al., 2021); (3) Cross-Modal Alignment, with (Chen et al., 2022b; Qin & Song, 2022) using attention mechanisms for text-visual feature correlation. Recent advancements (Gu et al., 2024; Tanida et al., 2023; Li et al., 2023c) focus on local visual feature learning and multi-level cross-modal alignment, showcasing the evolution of semantic matching in radiology report generation. However, these approaches mainly focus on radiology report generation, directly applying these approaches to ECG data is challenging because the patterns between medical images and ECG biomedical signals are significantly different.
Signal-to-Text Generation. Signal-to-Text (S2T) generation, predominantly speech-to-text, converts speech signals into text. Extensive research (Baevski et al., 2020; Chen et al., 2020; Liu et al., 2021a; Zheng et al., 2021; Hsu et al., 2021; Bapna et al., 2021) has extensively studied patterns in speech signals, analyzing both temporal and spectral aspects. Although the field of ECG report generation is largely unexplored, the use of speech signal research techniques in the creation of ECG reports presents significant challenges due to fundamental differences: (1) Speech signals for S2T tasks typically have a sampling frequency of up to 16,000 - 60,000 Hz (Lavry, 2012; Painter & Spanias, 2000; Byrne et al., 1994) and are monochannel, while ECG signals usually have a lower sampling frequency of around 500 Hz and encompass 12 channels. (2) Speech signals can be discretized into acoustic units corresponding to specific words (Défossez et al., 2022), but ECG signals lack such discrete units. Certain cardiac conditions, like bundle branch block, can obscure or distort the heartbeats patterns in ECG readings. (3) Speech signals and their transcriptions maintain strong temporal alignment, as each word correlates with a specific acoustic pattern, and the transcription follows the temporal sequence of speech (Zhang et al., 2023b). In contrast, ECG reports evaluate the cardiac patterns on a global level, assessing the entire ECG period rather than focusing on individual acoustic-like units in speech. These fundamental differences between speech-text pairs and ECG records, in both signal and text terms, highlight the need for S2T methods specifically designed for ECG data. Furthermore, the scarcity of research in the generation of ECG reports highlights an urgent need for comprehensive benchmarks using large-scale, publicly accessible data sets.
7 Conclusion and Future Work
In this work, we propose and demonstrate the efficacy of the Multimodal ECG Instruction Tuning (MEIT) framework. Particularly, we first introduce a novel method for creating instruction-following data for ECG reports, enabling the training of an ECG report generator, a multimodal LLM designed to generate ECG reports following human instructions. Furthermore, we propose an effective and efficient method for aligning ECG and report representations across multiple open-source LLMs, achieving notable performance on both the MIMIC-IV-ECG and PTB-XL datasets for diverse tasks. Importantly, this paper introduces the comprehensive benchmark for ECG instruction-following in ECG report generation. While this research primarily concentrates on ECG signals, it marks a preliminary step in biomedical signal instruction tuning. In future work, we plan to extend our framework to encompass additional medical signal domains, such as EEG signals, with the aspiration that our work spurs further advancements in developing more proficient medical-signal LLMs.
Broader Impact
Our MEIT framework first attempt to addresses automatic ECG report generation through ECG instruction tuning, establishing the first comprehensive benchmark for this process with mainstream LLMs. In this work, we mainly focus on generating reports from multimodal ECG instructions using LLMs. However, the generated results are not fully explainable or controllable, even though the generation procedure is transparent and trackable. This is because the underlying theory of LLMs remains largely unexplored, necessitating further investigation to ensure the quality and safety of the generated content. In the future, we aim to enable LLMs to utilize external, expert-verified knowledge databases, such as clinical protocols and medical textbooks, to enhance the explainability of the generated ECG reports.
Appendix A Hyper-parameters of ECG Instruction Tuning
Table 4: Hyper-parameters of ECG instruction tuning for all LLM backbones.
Hyperparameters
Mixed precision bf16
Instruction tuning epochs 5
LoRA alpha 64
LoRA rank 128
LoRA dropout 0.1
Total batch size 64
Gradient accumulation 2
Maximum sequence length 256
Learning rate 2e-5, 1e-4
Learning rate Optimizer AdamW
Schedule linear
Warm-up ratio 0.03
Weight decay 0.0
Table 5: ECG dimension of different language models.
Models ECG Dimension
GPT2-Medium 64
GPT2-Large 64
GPT-Neo 128
GPT-NeoX 96
GPT-J 256
BLOOM 128
OPT 128
LLaMA-1 128
Mistral 128
LLaMA-2 128
Mistral-Instruct 128
In this study, we concentrate on the implementation of the Low-Rank Adaptation (LoRA) (Hu et al., 2021) technique for efficient fine-tuning, specifically applied to ECG instruction tuning. As detailed in Table A provided, we utilize mixed precision at bf16 for enhanced computational efficiency. Our models undergo instruction tuning over 5 epochs, with LoRA parameters set at an alpha of 64 and a rank of 128, accompanied by a dropout rate of 0.1. The total batch size is 64 with a gradient accumulation factor of 2. The maximum sequence length is constrained to 256 tokens. Additionally, we adopt a learning rate with 2e-5 for GPT-NeoX, and 1e-4 for the other models, optimized using the AdamW algorithm. The learning rate follows a linear schedule with a warm-up ratio of 0.03. We set the weight decay to 0.0.
Moreover, as shown in Table A, we detail the ECG embedding dimensions for various language models, highlighting their approach to ECG data encoding. GPT2-Medium and GPT2-Large feature ECG dimensions of 64, while GPT-Neo, BLOOM, OPT, LLaMA-1, Mistral, LLaMA-2, and Mistral-Instruct use a dimension of 128. GPT-NeoX employs a dimension of 96, and GPT-J notably uses the largest dimension of 256. These dimensions, reflecting each model’s head dimension design, illustrate diverse strategies in ECG data processing across different models.
Appendix B Visualization of Generated ECG Report Samples
As illustrated in Figures 9,  10, and  11 we have visualized the report samples generated by LLaMA-1, LLaMA-2, and Mistral-Instruct. The samples are presented in blue font to highlight the key information that aligns with the ground truth. The visualization demonstrates that all three selected models are capable of capturing the essential patterns of ECG signals and generating accurate reports. This underscores the efficacy of our proposed MEIT framework, which proves to be adaptable to the majority of open-source LLMs. It effectively learns the correct clinical semantics of ECG signals, thereby enabling the generation of corresponding reports.
Figure 9: Reports generated by LLaMA-1 following ECG instruction Tuning.
Figure 10: Reports generated by LLaMA-2 following ECG instruction Tuning.
Figure 11: Reports generated by Mistral-Instruct following ECG instruction Tuning.

### v2

@@
 1 Introduction
-Diagnosing heart diseases is of paramount importance, requiring the analysis of diverse multimodal data sources, which includes electrocardiogram (ECG) recordings and clinical reports.
-Cardiologists interpret these ECG recordings to generate comprehensive ECG reports for heart disease diagnosis, which is a complex and time-consuming process.
-Recently, artificial intelligence (AI) models have been used to facilitate ECG data analysis by categorizing ECG data (Chen et al., 2022a; Hu et al., 2023; Liu et al., 2023a; Li et al., 2023b).
+Diagnosing heart diseases is of paramount importance, requiring the analysis of multimodal data sources, which includes electrocardiogram (ECG) recordings and clinical reports.
+Cardiologists read and interpret these ECG recordings to manually generate comprehensive ECG reports for heart disease diagnosis, which is a complex and time-consuming process.
+Recently, artificial intelligence (AI) models have been developed to facilitate ECG data analysis by categorizing ECG data (Chen et al., 2022a; Hu et al., 2023; Liu et al., 2023a; Li et al., 2023b).
 Despite these early efforts, the automatic generation of reports from ECG recordings remains an underexplored area.
@@
 To tackle these challenges, we introduce MEIT, a Multimodal ECG Instruction Tuning framework that extends the capabilities of LLMs in the cardiology context to generate ECG reports using ECG signals based on human instructions.
-Inspired by the versatility of LLMs (Achiam et al., 2023; Touvron et al., 2023b; Wan et al., 2023; Wang et al., 2024) in handling diverse language tasks simultaneously, we develop a specialized instruction tuning process for ECG report generation.
+Inspired by the versatility of LLMs (Achiam et al., 2023; Touvron et al., 2023b; Wan et al., 2023b; Wang et al., 2024) in handling diverse language tasks simultaneously, we develop a specialized instruction tuning process for ECG report generation.
 MEIT aligns human instructions with ECG signals, enabling LLMs to generate professional-grade reports and exhibit zero-shot report generation capabilities under domain transfer scenarios across various continents and data collection devices.
@@
 where $i$ is the current layer index.
-Different from gated-attention fusion in Flamingo (Alayrac et al., 2022) and Memorizing Transformer (Wu et al., 2022), or Q-former in BLIP-2 (Li et al., 2023a) that require the additional trainable parameters or design complex modules, we propose a simple yet effective concatenated-fusion method that directly injects the ECG embeddings with language context in the causal-attention (C-Attn) of the models.
+Different from gated-attention fusion in Flamingo (Alayrac et al., 2022), Memorizing Transformer (Wu et al., 2022) ,and G-MAP (Wan et al., 2022), or Q-former in BLIP-2 (Li et al., 2023a) that require the additional trainable parameters or design complex modules, we propose a simple yet effective concatenated-fusion method that directly injects the ECG embeddings with language context in the causal-attention (C-Attn) of the models.
 In our approach, each attention layer combines $\mathbf{H}_{e}$, generated from the ECG encoder and projector as a prefix condition, with $\mathbf{H}_{t}^{i}$, derived from the preceding layer.
@@
 Medical Report Generation.
-Medical Report Generation can be categorized into three main approaches: (1) Template Selection and Generation, where methods like HRGR (Li et al., 2018) focus on sentence creation or template retrieval, and CMAS (Jing et al., 2017) uses dual writers for differentiating regions; (2) Data Integration and Coherence, illustrated by PPKED’s (Liu et al., 2021b) knowledge merging for consistency, and CA’s contrasting of abnormal features (Ma et al., 2021); (3) Cross-Modal Alignment, with (Chen et al., 2022b; Qin & Song, 2022) using attention mechanisms for text-visual feature correlation.
+Different from medical vision-language alignment (Wan et al., 2023a), medical Report Generation can be categorized into three main approaches: (1) Template Selection and Generation, where methods like HRGR (Li et al., 2018) focus on sentence creation or template retrieval, and CMAS (Jing et al., 2017) uses dual writers for differentiating regions; (2) Data Integration and Coherence, illustrated by PPKED’s (Liu et al., 2021b) knowledge merging for consistency, and CA’s contrasting of abnormal features (Ma et al., 2021); (3) Cross-Modal Alignment, with (Chen et al., 2022b; Qin & Song, 2022) using attention mechanisms for text-visual feature correlation.
 Recent advancements (Gu et al., 2024; Tanida et al., 2023; Li et al., 2023c) focus on local visual feature learning and multi-level cross-modal alignment, showcasing the evolution of semantic matching in radiology report generation.

### v3

@@
-Electrocardiogram Instruction Tuning for Report Generation
+MEIT: Multi-Modal Electrocardiogram Instruction Tuning on Large Language Models for Report Generation
 Abstract
-Electrocardiogram (ECG) serves as the primary non-invasive diagnostic tool for cardiac conditions monitoring, are crucial in assisting clinicians.
-Recent studies have concentrated on classifying cardiac conditions using ECG data but have overlooked ECG report generation, which is not only time-consuming but also requires clinical expertise.
+Electrocardiogram (ECG) is the primary non-invasive diagnostic tool for monitoring cardiac conditions and is crucial in assisting clinicians.
+Recent studies have concentrated on classifying cardiac conditions using ECG data but have overlooked ECG report generation, which is time-consuming and requires clinical expertise.
 To automate ECG report generation and ensure its versatility, we propose the Multimodal ECG Instruction Tuning (MEIT) framework, the first attempt to tackle ECG report generation with LLMs and multimodal instructions.
 To facilitate future research, we establish a benchmark to evaluate MEIT with various LLMs backbones across two large-scale ECG datasets.
-Our approach uniquely aligns the representations of the ECG signal and the report, and we conduct extensive experiments to benchmark MEIT with nine open source LLMs, using more than 800,000 ECG reports.
+Our approach uniquely aligns the representations of the ECG signal and the report, and we conduct extensive experiments to benchmark MEIT with nine open-source LLMs using more than 800,000 ECG reports.
 MEIT’s results underscore the superior performance of instruction-tuned LLMs, showcasing their proficiency in quality report generation, zero-shot capabilities, and resilience to signal perturbation.
-These findings emphasize the efficacy of our MEIT [^1] All data and code will be available upon acceptance.
-framework and its potential for real-world clinical application.
-Machine Learning, ICML
-[…]
-[…]
-${}^{\text{1}}$he Ohio State University ${}^{\text{2}}$niversity of North Carolina at Chapel Hill ${}^{\text{3}}$mperial College London
-${}^{\text{4}}$ King’s College London  ${}^{\text{5}}$ ${}^{\text{6}}$
-$*{}^{*}$orrespondence to: Zhongwei Wan<wan.512@osu.edu>Mi Zhang<mizhang.1@osu.edu>
-${}^{\text{\textdagger}}$ore contribution  ${}^{\ddagger}$ore advising
+These findings emphasize the efficacy of our MEIT framework and its potential for real-world clinical application.
 1 Introduction
-Diagnosing heart diseases is of paramount importance, requiring the analysis of multimodal data sources, which includes electrocardiogram (ECG) recordings and clinical reports.
+Diagnosing heart diseases is of paramount importance, requiring the analysis of multimodal data sources, which include electrocardiogram (ECG) recordings and clinical reports.
 Cardiologists read and interpret these ECG recordings to manually generate comprehensive ECG reports for heart disease diagnosis, which is a complex and time-consuming process.
-Recently, artificial intelligence (AI) models have been developed to facilitate ECG data analysis by categorizing ECG data (Chen et al., 2022a; Hu et al., 2023; Liu et al., 2023a; Li et al., 2023b).
-Despite these early efforts, the automatic generation of reports from ECG recordings remains an underexplored area.
-Unlike other AI-empowered medical report generation applications (e.g., radiology report), the primary challenges for ECG report generation stem from the distinct nature of ECG content.
+Recently, AI models have been developed to facilitate ECG data analysis for the task of classification [1, 2, 3, 4].
+Despite these efforts, the automatic generation of reports from ECG recordings still needs to be explored.
+Unlike other AI-empowered medical report generation applications (e.g., radiology reports), the primary challenge for ECG report generation stems from the distinct nature of ECG content.
 ECG reports, often comprising brief phrases that summarize signal patterns, contrast with detailed anatomical descriptions in radiology reports.
-This difference in content and semantic interpretation between imaging and signal data complicates the direct application of radiology-focused methods to ECG reports.
+The difference in the content and semantic interpretation between imaging and ECG data complicates the direct application of radiology-focused methods to ECG reports.
 Furthermore, there is still a lack of comprehensive benchmarks for evaluating the performance of ECG report generation.
-To tackle these challenges, we introduce MEIT, a Multimodal ECG Instruction Tuning framework that extends the capabilities of LLMs in the cardiology context to generate ECG reports using ECG signals based on human instructions.
-Inspired by the versatility of LLMs (Achiam et al., 2023; Touvron et al., 2023b; Wan et al., 2023b; Wang et al., 2024) in handling diverse language tasks simultaneously, we develop a specialized instruction tuning process for ECG report generation.
-MEIT aligns human instructions with ECG signals, enabling LLMs to generate professional-grade reports and exhibit zero-shot report generation capabilities under domain transfer scenarios across various continents and data collection devices.
+To tackle these challenges, we introduce MEIT, a Multimodal ECG Instruction Tuning framework that extends the capabilities of LLMs in the cardiology context to generate ECG reports using ECG recordings and human instructions.
+Inspired by the versatility of LLMs [5, 6, 7, 8, 9] in handling diverse language tasks simultaneously, we develop a specialized instruction tuning process for ECG report generation.
+MEIT aligns human instructions with ECG recordings, enabling LLMs to generate professional-grade reports and exhibit zero-shot report generation capabilities under domain transfer scenarios across various continents and data collection devices.
 Specifically, leveraging publicly available ECG datasets, we construct a multimodal instruction dataset including ECG records, human instructions, and paired reports.
-Then, we propose an effective and efficient attention-based fusion method to integrate ECG and text representations in the latent space, enabling LLMs to understand signals for report generation without introducing additional training parameters in the attention layer.
-In addition to the ECG report generation approach within MEIT, we introduce a comprehensive benchmark for ECG report evaluation, utilizing two datasets with 20K and 800K ECG-report pairs, respectively, across three evaluation tasks: report generation quality, zero-shot learning across datasets, and robustness analysis in the face of ECG signal perturbation.
+Then, we propose an effective and efficient attention-based fusion method to integrate ECG and text representations in the latent space.
+This enables LLMs to understand ECG signals for report generation without introducing additional training parameters in the attention layer.
+In addition to the ECG report generation approach, we introduce a comprehensive benchmark for ECG report generation evaluation, utilizing two datasets with 20K and 800K ECG-report pairs, respectively, across three evaluation tasks: report generation quality, zero-shot learning across datasets, and robustness analysis in the face of ECG signal perturbation.
 Utilizing the ECG report evaluation benchmark, we assess the proposed approach across nine open-source LLMs.
@@
 To summarize, our primary contribution is the MELT framework, a novel approach to automating ECG report generation and evaluation based on LLMs.
-This framework incorporates a lightweight, attention-based fusion module effective across various LLM models.
-Furthermore, we establish a new benchmark for ECG report generation, which contains three aspects of the evaluation tasks.
-Our evaluations showcase the enhanced capabilities of instruction-tuned LLMs in generating ECG reports, highlighting the transferability in zero-shot test and robustness against data perturbations.
-The MEIT framework paves the way for future advances in automated ECG report generation and methodological innovations in integrating biomedical signals into LLMs.
+This framework incorporates a lightweight, attention-based fusion module across various LLM models.
+Furthermore, we design a new benchmark for ECG report generation, which contains three evaluation tasks.
+Our evaluations showcase the enhanced capabilities of instruction-tuned LLMs in generating ECG reports, highlighting the transferability in zero-shot tests and robustness against data perturbations.
+MEIT paves the way for future advances in automated ECG report generation and methodological innovations in integrating biomedical signals into LLMs.
 2 Preliminaries
 Electrocardiogram (ECG) measures the electrical activity of an individual’s heart over time.
-An ECG recording typically contains a 12-lead multivariate time series, which act as a 12-dimensional sequence of embeddings.
+An ECG recording typically contains a 12-lead multivariate time series, which acts as a 12-dimensional sequence of embeddings.
 The ECG signal offers a comprehensive view, encompassing both spatial and temporal aspects of cardiac function.
-ECG leads can be categorized into 6 limb leads (i.e., I, II, III, aVR, aVL, and aVF) to monitor arms and legs, providing frontal plane views, and 6 precordial leads (i.e., V1, V2, V3, V4, V5, and V6) to monitor chest, showing horizontal plane views.
-We denote an ECG recording as $\mathbf{X}_{e}\in\mathcal{R}^{M\times T}$, where $M$ represents the numbers of leads, $T$ is signal length.
+ECG leads can be categorized into six limb leads (i.e., I, II, III, aVR, aVL, and aVF) to monitor arms and legs, providing frontal plane views, and six precordial leads (i.e., V1, V2, V3, V4, V5, and V6) to monitor chest, showing horizontal plane views.
+We denote an ECG recording as $\mathbf{X}_{e}\in\mathcal{R}^{M\times T}$, where $M$ represents the number of leads, and $T$ is signal length.
 Each ECG recording is associated with an ECG report $\mathbf{X}_{t}$ for description.
 Thus, we denote each ECG pair as $\{\mathbf{X}_{e},\mathbf{X}_{t}\}$.
-More visualization details can be found in the Appendix.
+More details on visualization can be found in the appendix A.5.
 Figure 1: Illustration of the proposed MEIT framework, including ECG instruction data curation, instruction tuning on auto-regressive objective, and inference to get the generated report.
@@
 We then delve into the architecture of the multimodal LLM designed for ECG report generation.
-Lastly, we describe the approach used to generate ECG reports with instruction tuning.
+Lastly, we describe the approach to generating ECG reports with instruction tuning.
 3.1 ECG Instruction Data Curation
 Given an ECG signal $\mathbf{X}_{e}$, our goal during inference is to generate an ECG report using an instruction prompt.
-For instance, the prompt might be, “Given the ECG signal embeddings, please help me generate an accurate description for this ECG signal embeddings: ”.
+For instance, the prompt can be “Given the ECG signal embeddings, please help me generate an accurate description for this ECG signal embeddings: ".
 To achieve this goal, during the training phase, we aim to create instruction tuning data to generate a response $\mathbf{\hat{X}}_{t}$ that aligns semantically with the ground truth $\mathbf{X}_{t}$.
 In addition, since we cannot predict the exact instruction prompt that users will use, we need to ensure that our report generation process is robust enough to handle different prompts.
-To address this challenge, we first manually design some prompt samples, then utilize GPT-4 (Achiam et al., 2023) to generate a set of prompts by rephrasing, as shown in Figure 1.
-Then, we randomly select one instruction prompt $\mathbf{X}_{p}$ from the prompt set, and create a general instruction-following template: <|user|>: $\mathbf{X}_{p}$, $\mathbf{X}_{e}$}<|assistant|>: $\mathbf{X}_{t}$} </s>, where <|user|> and <|assistant|> are added special tokens for tokenizer, </s> is a stop sign for each response.
-This approach ensures that the generated response not only conveys the same meaning as the ground truth but also remains adaptable to different instruction prompts.
-Following this strategy, we construct the ECG instruction data using a MIMIC-IV-ECG (Gow et al., ) dataset that contains 80K annotated data, and a 20K dataset PTB-XL (Wagner et al., 2020).
+To address this challenge, we manually design some prompt samples, then utilize GPT-4 [5] to generate a set of prompts by rephrasing, as shown in Figure 2.
+Then, we randomly select one instruction prompt $\mathbf{X}_{p}$ from the prompt set and create a general instruction-following template: <|user|>: $\mathbf{X}_{p}$, $\mathbf{X}_{e}$}<|assistant|>: $\mathbf{X}_{t}$} </s>, where <|user|> and <|assistant|> are added special tokens for tokenizer, </s> is a stop sign for each response.
+This approach ensures that the generated response conveys the same meaning as the ground truth and remains adaptable to different instruction prompts.
+Following this strategy, we construct the ECG instruction data using a MIMIC-IV-ECG [10] dataset that contains 80K annotated data and a 20K dataset PTB-XL [11].
+The ECG instruction data samples are shown in Appendix A.5.
 3.2 ECG Report Generation Model
-In MEIT framework, the multimodal ECG report generation model aims to decode ECG signals end-to-end and accurately generate ECG reports.
+In MEIT, the multimodal ECG report generation model decodes ECG signals end-to-end to generate ECG reports.
 The architecture is illustrated in Figure 2.
-Specifically, the report generation model aims to encode the ECG-signal $\mathbf{X}_{e}\in\mathcal{R}^{M\times T}$ into latent embeddings and integrate them with the language embeddings with modality alignment, and then autoregressively generate the ECG report.
-In the following, we detail each component of the report generation model.
+Specifically, the report generation model directly encodes an entire ECG-signal $\mathbf{X}_{e}\in\mathcal{R}^{M\times T}$ into latent embeddings and integrate it with the language embeddings with modality alignment, and then autoregressively generate the ECG report.
+Next, we detail each component of the report generation model.
 ECG Encoder.
-Since ECG signal is of high resolution in the temporal domain, it is vital to efficiently extract temporal features per lead before interaction with semantic embeddings inside the LLM backbone.
-Our default ECG encoder $\mathcal{F}_{e}(\cdot)$ consists of some temporal convolution blocks to encode each ECG signal into embeddings.
-Specifically, each temporal convolution block is composed of several 1-D convolution layers, batch normalization layers, and ReLU activation layers followed by average pooling.
-To align the output dimension with the head dimension of LLM backbone $\mathcal{F}_{l}(\cdot)$, we employ a non-linear up-to-down projection layer $\mathcal{P}_{e}(\cdot)$ to obtain the ECG embeddings:
+Since the ECG signal is of high resolution in the temporal domain, it is vital to efficiently extract temporal features per lead before interaction with semantic embeddings inside the LLM backbone.
+Our default ECG encoder $\mathcal{F}_{e}(\cdot)$ consists of temporal convolution blocks to encode each ECG signal into embeddings.
+Specifically, each temporal convolution block comprises several 1-D convolution layers, batch normalization layers, and ReLU activation layers, followed by average pooling.
+To align the output dimension with the head dimension of LLM backbone $\mathcal{F}_{l}(\cdot)$, we employ a non-linear projection layer $\mathcal{P}_{e}(\cdot)$ to obtain the ECG embeddings:
 $\mathbf{H}_{e}=\mathcal{P}_{e}\left(\mathcal{F}_{e}\left(\mathbf{X}_{e}\right)\right),$ (1)
@@
 Note that our default ECG encoder is lightweight and is able to learn temporal patterns of signals without a long training period.
+More details about ECG Encoder are illustrated in Appendix A.3.
 ECG Modality Alignment.
-To guide the LLMs in aligning ECG signal data with corresponding textual outputs, we introduce an ECG modality alignment strategy.
+We introduce an ECG modality alignment strategy to guide the LLMs in aligning ECG signal data with corresponding textual outputs.
 This approach is detailed in Figure 2.
-Specifically, given the ECG embeddings $\mathbf{Z}_{e}$, the alignment strategy aims to incorporate $\mathbf{Z}_{e}$ with the current hidden state $\mathbf{H}_{t}^{i}$ generated from previous $i-1$-th layer of the LLM backbone $\mathcal{F}_{l}\left(\cdot\right)$ for next-token prediction task.
+Specifically, given the ECG embeddings $\mathbf{H}_{e}$, the alignment strategy incorporates $\mathbf{H}_{e}$ with the current hidden state $\mathbf{H}_{t}^{i}$ generated from previous $i-1$-th layer of the LLM backbone $\mathcal{F}_{l}\left(\cdot\right)$ for next-token prediction task.
 Here $\mathbf{H}_{t}^{i}$ is defined as:
@@
 where $i$ is the current layer index.
-Different from gated-attention fusion in Flamingo (Alayrac et al., 2022), Memorizing Transformer (Wu et al., 2022) ,and G-MAP (Wan et al., 2022), or Q-former in BLIP-2 (Li et al., 2023a) that require the additional trainable parameters or design complex modules, we propose a simple yet effective concatenated-fusion method that directly injects the ECG embeddings with language context in the causal-attention (C-Attn) of the models.
+Different from gated-attention fusion in Flamingo [12], Memorizing Transformer [13] , and G-MAP [14], or Q-former in BLIP-2 [15] that require the additional trainable parameters and design complex modules for two-stage training, we propose a simple yet effective concatenated-fusion method that directly injects the ECG embeddings with language context in the causal-attention (C-Attn) of the models.
 In our approach, each attention layer combines $\mathbf{H}_{e}$, generated from the ECG encoder and projector as a prefix condition, with $\mathbf{H}_{t}^{i}$, derived from the preceding layer.
@@
 $\operatorname{C-Attn}\left(\mathbf{H}_{e},\mathbf{H}_{t}^{i}\right)=\left[\operatorname{head}_{1},\ldots,\operatorname{head}_{k}\right]\mathbf{W}_{o},$ (3)
-here, $k$ represents the number of attention heads, and $\mathbf{W}_{o}$, a matrix in $\mathcal{R}^{kD_{h}\times D_{m}}$, serves as the projection matrix with $D_{m}$ denoting the hidden size of the LLM backbone.
+where $k$ represents the number of attention heads, and $\mathbf{W}_{o}$, a matrix in $\mathcal{R}^{kD_{h}\times D_{m}}$, serves as the projection matrix with $D_{m}$ denoting the hidden size of the LLM backbone.
 We replicate $\mathbf{H}_{e}$ for each head $k$ times, merging the ECG and language features in the sequence dimension.
@@
 Our model’s design allows for the efficient fusion of two modalities through causal attention, facilitating conditional generation without the need for additional parameter updates to align the ECG modality.
+More comparisons about ECG modality alignment and other fusion approaches are illustrated in Appendix LABEL:appendix:_ECG_align_ablation.
 3.3 ECG Instruction Tuning
@@
 During instruction tuning, we compute autoregressive loss only on tokens after response tokens <assistant>, and use label loss masking to finetune the model, where we mask all tokens belonging to $\mathbf{X}_{p}$ and $\mathbf{X}_{e}$.
-To save computational resources and accelerate the convergence of instruction tuning, we use the LoRA (Hu et al., 2021) adapters for all linear layers of the LLM backbone $\mathcal{F}_{l}$ and freeze its backbone.
+To save computational resources and accelerate the convergence of instruction tuning, we use LoRA [16] adapters for all linear layers of the LLM backbone $\mathcal{F}_{l}$ and freeze its backbone.
 Subsequently, given a sequence of ECG instruction data, we compute the probability of the target response $\mathbf{X_{t}}$ as an autoregressive function:
@@
 4 Benchmarking ECG Report Generation
+In this section, we create a benchmark to evaluate the ECG report generation approach.
+Our benchmark includes two datasets (PTB-XL and MIMIC-IV-ECG) along with three ECG-related tasks, including (1) the quality of ECG report generation, (2) the zero-shot ability under domain transfer, (3) the robustness of the model against perturbed signals.
+4.1 Datasets
+PTB-XL.
+The PTB-XL dataset [11] comprises 21,837 clinical 12-lead ECGs, each lasting 10 seconds, from 18,885 patients.
+Each ECG is accompanied by a corresponding report.
+The data includes raw ECG signals and their reports, with all recordings sampled at 500Hz over 10 seconds.
+We divided this dataset into training, validation, and testing subsets in a 70%:10%:20% ratio, respectively.
+The human experts double-check all samples in the test set to ensure data quality.
+As mentioned in Sec 3.1, we reformulate the dataset into the instruction data format.
+MIMIC-IV-ECG.
+The MIMIC-IV-ECG dataset [10] is currently the largest publicly accessible ECG dataset, containing 800,035 paired samples from 161,352 unique subjects.
+Similar to PTB-XL, each sample in this dataset includes a raw ECG signal and its corresponding report, with all recordings sampled at 500Hz for 10 seconds.
+The division of this dataset into training, validation, and testing subsets is 80%:10%:10% ratio.
+Likewise, we reconstruct this dataset into an ECG instruction data template.
+4.2 Evaluation Metrics
+We evaluate models using nine metrics: BLEU 1-4 [17], METEOR [18], ROUGE 1-2 and L [19], CIDEr-D [20], and BERTScore [21].
+BLEU and METEOR assess machine translation quality, focusing on accuracy and fluency.
+ROUGE-L measures sentence fluency and structure, while ROUGE-1 and ROUGE-2 examine uni-gram and bi-gram overlaps.
+CIDEr-D evaluates the relevance and uniqueness of generated ECG reports against a candidate set, and BERTScore assesses semantic similarity to the ground truth, ensuring content accuracy.
+4.3 Tasks
+We evaluate the performance of ECG report generation on the following three tasks:
+Quality Report Generation.
+This task aims to assess report quality after ECG instruction tuning using 10% of MIMIC-IV-ECG and PTB-XL datasets as the test set.
+The evaluation examines how closely generated reports match the original’s structure and meaning, considering various instructions and ECG inputs.
+We analyze metrics like BLEU-1 to 4, METEOR, ROUGE 1, 2, L, CIDEr-D, and BERTScore.
+Zero-shot Generalizability.
+To explore the generalizability of LLMs in domain transfer scenarios following ECG instruction tuning, we trained the models on 70% of the instruction data from MIMIC-IV-ECG.
+Following this, we evaluated the models’ zero-shot capabilities on the PTB-XL test set.
+It’s important to note that the PTB-XL and MIMIC-IV-ECG datasets originate from different continents—Europe and the United States, respectively—utilizing varied devices and from distinct hospitals, across different time periods.
+Therefore, we consider these datasets to represent two separate domains.
+This distinction allows us to use the PTB-XL dataset to gauge our model’s performance in zero-shot domain transfer effectively.
+We used the metrics BLEU-4, METEOR, ROUGE-L, and CIDEr-D because of limited space and calculated their average for model evaluation.
+Signal Perturbation Robustness.
+In real-world clinical settings, ECG signals often contain some degree of noise.
+To evaluate the robustness of MEIT against such noisy interference, we selected 10% of the ECG samples from the MIMIC-IV-ECG test dataset.
+We then added Gaussian noise to these samples during the models’ instruction-based inference process.
+For this evaluation, we used BLEU-4, METEOR, ROUGE-L, and CIDEr-D as metrics.
+5 Experiments and Evaluation
+5.1 Experimental Settings
+In this section, we evaluate and benchmark nine open-source decoder-only LLMs using the constructed ECG report generation benchmark.
+Additionally, we offer a comprehensive analysis of scalability and instruction tuning and present case studies showcasing the generated reports.
+Models.
+We use nine LLMs based on the peft [^1] https://github.com/huggingface/peft library, which directly supports LoRA [16] to construct the multimodal ECG report generation model described in Section 3.2.
+These models include GPT-Neo [22], GPT-NeoX [23], GPT-J [24], BLOOM [25], OPT [26], LLaMA-1 [27], LLaMA-2 [6], Mistral [28], and Mistral-Instruct [^2] https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.1, along with
 Table 1: Natural language generation metric on MIMIC-IV-ECG.
@@
 Mistral-Instruct† 7B 0.501 0.481 0.457 0.425 0.664 0.592 0.700 0.641 4.01
+two relatively small pre-trained language models (GPT2-Medium and GPT-Large [29]) as fundamental baselines.
+Implementation Details. Our study utilized PyTorch 2.1, transformers [30], and accelerated on A100
 Table 3: Semantic similarity between the generated ECG reports and ground truths is measured using BERTScore, denoted as P for Precision, R for Recall, and F-1 for the F-1 Score.
@@
 Mistral-Instruct† 0.773 0.722 0.747 0.730 0.661 0.694
-In this section, we create a benchmark to evaluate the ECG report generation approach.
-Our benchmark includes two datasets (PTB-XL and MIMIC-IV-ECG) along with three ECG-related tasks, including (1) the quality of ECG report generation; (2) the zero-shot ability under domain transfer; (3) the robustness of model against perturbed signals.
-4.1 Datasets
-PTB-XL.
-The PTB-XL dataset (Wagner et al., 2020) comprises 21,837 clinical 12-lead ECGs, each lasting 10 seconds, from 18,885 patients.
-Each ECG is accompanied by a corresponding report.
-The data includes raw ECG signals and their reports, with all recordings sampled at 500Hz over 10 seconds.
-We divided this dataset into training, validation, and testing subsets in a 70%:10%:20% ratio, respectively.
-The human experts double-check all samples in the test set to ensure data quality.
-We reformulate the dataset into instruction data format as mentioned in Sec 2.
-MIMIC-IV-ECG.
-The MIMIC-IV-ECG dataset (Gow et al., ) is currently the largest publicly accessible ECG dataset, containing 800,035 paired samples from 161,352 unique subjects.
-Similar to PTB-XL, each sample in this dataset includes a raw ECG signal and its corresponding report, with all recordings sampled at 500Hz for 10 seconds.
-The division of this dataset into training, validation, and testing subsets is 80%:10%:10% ratio.
-Likewise, we reconstruct this dataset into an ECG instruction data template.
-4.2 Evaluation Metrics
-We evaluate models using nine metrics: BLEU 1-4 (Papineni et al., 2002), METEOR (Banerjee & Lavie, 2005), ROUGE 1-2 and L (Lin, 2004), CIDEr-D (Vedantam et al., 2015), and BERTScore (Zhang et al., 2019).
-BLEU and METEOR assess the quality of machine translation, focusing on accuracy and fluency.
-ROUGE-L measures sentence fluency and structure, while ROUGE-1 and ROUGE-2 examine uni-gram and bi-gram overlaps.
-CIDEr-D evaluates the relevance and uniqueness of generated ECG reports against a candidate set, and BERTScore assesses semantic similarity to the ground truth, ensuring content accuracy.
-4.3 Tasks
-We evaluate the performance of ECG report generation on the following three tasks:
-Quality Report Generation.
-The goal of this task is to evaluate the quality of report generation following ECG instruction tuning.
-We utilize 10% of the MIMIC-IV-ECG and PTB-XL datasets, designated for instruction tuning, as the test set.
-This evaluation focuses on the syntactic and semantic similarities between the generated ECG report and the ground truth, based on diverse instructions and input ECG signals.
-For task assessment, we employ BLEU-1 to 4, METEOR, ROUGE 1, 2, and L, CIDEr-D, and BERTScore.
-Zero-shot Generalizability.
-To explore the generalizability of LLMs in domain transfer scenarios following ECG instruction tuning, we trained the models on 70% of the instruction data from MIMIC-IV-ECG.
-We then assessed the zero-shot capabilities of these models on the PTB-XL test set, which contains high-quality ground truth evaluated by clinical experts.
-For model evaluation, we used the metrics BLEU-3, BLEU-4, METEOR, and ROUGE-L and calculated their average.
-Signal Perturbation Robustness.
-In real-world clinical settings, ECG signals often contain some degree of noise.
-To evaluate the robustness of MEIT against such noisy interference, we selected 10% of the ECG samples from the MIMIC-IV-ECG test dataset.
-We then added Gaussian noise to these samples during the models’ instruction-based inference process.
-For this evaluation, we used BLEU-4, METEOR, ROUGE-L, and CIDEr-D as metrics.
-5 Experiments and Evaluation
-In this section, we evaluate and benchmark nine open-source decoder-only LLMs using the constructed ECG report generation benchmark.
-Additionally, we offer a comprehensive analysis of scalability and instruction tuning and present case studies showcasing the generated reports.
-5.1 Experimental Settings
-Models.
-We use nine LLMs based on the peft [^2] https://github.com/huggingface/peft library, which directly supports the LoRA (Hu et al., 2021) module, to construct the multimodal ECG report generation model described in Section 3.2.
-These models include GPT-Neo (Black et al., 2021), GPT-NeoX (Black et al., 2022), GPT-J (Wang & Komatsuzaki, 2021), BLOOM (Workshop et al., 2022), OPT (Zhang et al., 2022), LLaMA-1 (Touvron et al., 2023a), LLaMA-2 (Touvron et al., 2023b), Mistral (Jiang et al., 2023), and Mistral-Instruct [^3] https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.1, along with two relatively small pre-trained language models (GPT2-Medium and GPT-Large (Radford et al., 2019)) as fundamental baselines.
-Implementation Details.
-Our experiment is implemented using PyTorch 2.1, the latest transformers (Wolf et al., 2020), and the accelerate (Gugger et al., 2022) library on 4 $\times$ A100 GPUs.
-We utilize the official weights for all LLMs from Hugging Face (Wolf et al., 2019) for evaluation, with sizes ranging from 2.7 billion to 70 billion parameters.
-For models exceeding 10 billion parameters, we use DeepSpeed [^4] https://github.com/microsoft/DeepSpeed for training.
-All models are trained on both MIMIC-IV-ECG and PTB-XL instruction datasets for 5 epochs with a learning rate of 2e-5 and a batch size of 64.
-We adopt a linear optimizer scheduler with a warm-up ratio of 0.03.
-The best performance of each model is determined by training on the training set, selecting from the validation set, and evaluating on the test set.
-More details about instruction tuning can be found in the appendix.
+GPUs with LLMs from Hugging Face [31] ranging from 2.7 to 70 billion parameters.
+For larger models, we used DeepSpeed [^3] https://github.com/microsoft/DeepSpeed.
+The training covered 5 epochs on MIMIC-IV-ECG and PTB-XL with a 2e-5 learning rate and 64 batch size, employing a linear optimizer with a 0.03 warm-up ratio.
+For text preprocessing, we initially remove all instances of the ‘nan’ string and sentences that consist solely of numerical values.
+Subsequently, we discard any samples whose reports contain fewer than 5 tokens.
+For, ECG encoder, we adopt random initialization.
+Additionally, the default number of generated prompts from GPT-4 is 256, more training, visualization details about ECG instruction tuning are illustrated in Appendix A.2, Section  5.3, and Appendix A.4.
 5.2 Benchmark Task Evaluation
@@
 Performance on MIMIC-I V-ECG.
-Table 1 and 3 present the results of various types of language encoders $\mathcal{F}_{l}(\cdot)$ on MIMIC-IV-ECG.
-The results show that all LLMs exhibit superior performance compared to smaller language models (SLMs), such as GPT2-Medium and GPT2-Large, across all evaluation metrics.
+Table 3 and 1 present the results of various types of language encoders $\mathcal{F}_{l}(\cdot)$ on MIMIC-IV-ECG.
+The results show that all LLMs perform better than smaller language models (SLMs), such as GPT2-Medium and GPT2-Large, across all evaluation metrics.
 Notably, from GPT-Neo to Mistral-Instruct, LLM-based backbones achieve a significant margin over SLMs in all metrics.
 For instance, compared to GPT2-Large, the METEOR score increases in the range of 0.132 to 0.18 from GPT-Neo to LLaMA-2, and Mistral-Instruct outperforms GPT2-Large with an improvement of 0.18 in the ROUGE-L score and 0.134 in the F-1 of BERTScore.
+Figure 3: Zero-shot performance on PTB-XL dataset. “IT” denotes instruction tuning.
 The observed performance underscores the adeptness of LLMs in generalizing from signal data, showcasing enhanced proficiency in aligning ECG signal representations with corresponding textual information.
-This highlights the significant potential of LLMs in the domain of medical signal-to-text generation.
-Particularly, LLaMA-2 and Mistral-Instruct surpass their counterparts in the majority of evaluative metrics, suggesting that models pre-tuned with general instructions are more adept at learning ECG-text alignment.
-Figure 3: Zero-shot performance on PTB-XL dataset. “IT” denotes instruction tuning.
-Figure 4: Signal perturbation robustness analysis on various LLMs
+This highlights the significant potential of LLMs in medical signal-to-text generation.
+Particularly, LLaMA-2 and Mistral-Instruct surpass their counterparts in most evaluative metrics, suggesting that models pre-tuned with general instructions are more adept at learning ECG-text alignment.
 Performance on PTB-XL.
@@
 Specifically, LLaMA-2 surpasses GPT2-Large by 0.134 in the BLEU-3 metric, while LLaMA-1 achieves a 0.103 improvement in the METEOR score.
-The overall experimental results reveal that Mistral-Instruct and LLaMA-2 are consistently the top two performers across most metrics, because of their strong general instruction-following capabilities.
-Consequently, we select LLaMA-2 and Mistral-Instruct as the default architectures for our framework.
+The overall experimental results reveal that Mistral-Instruct and LLaMA-2 are consistently the top two performers across most metrics because of their strong general instruction-following capabilities.
 5.2.2 Zero-shot Evaluation in Domain Transfer.
@@
 PTB-XL IT represents training on the PTB-XL train set and then evaluated on the PTB-XL test set.
-Notably, although Zero-shot IT shows a slight degradation compared to PTB-XL IT, the results still indicate a variance in the models’ ability to generalize to an unseen dataset with instruction tuning (IT), compared to Zero-shot W/O IT.
+Notably, although Zero-shot IT shows a slight degradation compared to PTB-XL IT, the results still indicate a variance in the model’s ability to generalize to an unseen dataset with instruction tuning (IT), compared to Zero-shot W/O IT.
 The involvement of ECG instruction tuning on MIMIC-IV-ECG enables the models to achieve superior zero-shot performance on the unseen PTB-XL dataset, indicating the necessity of instruction tuning in enhancing the models’ zero-shot ability on unseen datasets in ECG report generation.
-Figure 5: Ablation Study of ECG Instruction Tuning on MIMIC-IV-ECG Dataset.
 5.2.3 Robust Analysis with Perturbed ECG Signal.
-In a noise stress evaluation (Wang et al., 2019), we added Gaussian noise to ECG signals at signal-to-noise ratios (SNRs) of 0.05, 0.1, 0.15, and 0.2 during testing to assess model robustness.
-Our experiments utilized four LLM architectures: BLOOM, OPT, LLaMA-1, and Mistral, each trained on clean ECG signals from the MIMIC-IV-ECG training set and tested on corresponding noise-added signals from its test set.
-The results, illustrated in Figure 4, show a performance decline in all LLMs as SNR increased, highlighting the significant interference of ECG noise.
+In a noise stress evaluation [32], we added Gaussian noise to ECG signals at signal-to-noise ratios (SNRs) of 0.05, 0.1, 0.15, and 0.2 during testing to assess model robustness.
+Our experiments utilized four LLM architectures: BLOOM, OPT, LLaMA-1, and Mistral, each trained on clean ECG signals from the MIMIC-IV-ECG
+Figure 4: Signal perturbation robustness analysis on various LLMs
+training set and tested on corresponding noise-added signals from its test set.
+The results, illustrated in Figure 4, show a performance decline in all LLMs as SNR decreased, highlighting the significant interference of ECG noise.
 Furthermore, as shown in Table 1, Mistral also excelled in tests on noise-free datasets, suggesting a synergistic effect between clean and noisy test sets.
 The results demonstrate Mistral’s strong resistance to perturbations.
-Even with more severe noise, it maintained robustness in terms of ROUGE-L and METEOR metrics.
+Even with more severe noise, it maintained robustness regarding ROUGE-L and METEOR metrics.
 Developing an even more robust framework is a goal for future research.
+Table 4: Performance comparison of the proposed concatenated-fusion method and other mainstream fusion variants.
+We evaluate these methods on the MIMIC-IV-ECG dataset, using BLEU-4, METEOR, ROUGE-L, and CIDEr-D metrics.
+We take LLaMA-1 7B as the LLM backbone here.
+heavy teal color indicates the highest results.
+Framework Method BLEU-4 METEOR ROUGE-L CIDEr-D
+LLaVA Straightforward input 0.529 0.737 0.712 4.99
+Flamingo Trainable cross-attention 0.527 0.768 0.715 5.11
+MEIT Concatenated-fusion 0.543 0.761 0.724 5.26
 5.3 Analysis
-5.3.1 Instruction Tuning Visualization.
+Instruction Tuning Visualization.
 Figure 5 compares the convergence curves of the instruction tuning loss and the METEOR score between GPT-Neo (2.7B), BLOOM (7B), OPT (6.7B), and LLaMA-2 (7B) on the MIMIC-IV-ECG train and validation datasets.
-We observe that larger models with more parameters can converge to a smaller loss and achieve higher performance on the METEOR score.
+We observe that larger models with more parameters can converge to a more minor loss and achieve higher performance on the METEOR score.
 Notably, an increase in model size correlates with higher performance and lower loss, suggesting that larger models have the potential for better performance.
+Analysis of ECG Modality Alignment.
+To study the effectiveness of our proposed concatenated-fusion method for ECG modality alignment, we compare it with other fusion approaches such as direct input in LLaVA [33] and additional trainable cross-attention layer in Flamingo [12].
+For straightforward input, we follow the design of LLaVA by directly concatenating the ECG encoder’s output embeddings with the sentence’s embeddings before inputting them into the LLM backbones.
+For the second comparison method, we follow Flamingo by adding a trainable cross-attention layer within the attention block.
+From Table 4, we observe that the Concatenated-fusion method outperforms the trainable cross-attention method of Flamingo in most metrics and is consistently superior to the Straightforward input method of LLaVA.
+Consequently, the concatenated fusion is more effective for the LLM backbone’s alignment with fine-grained ECG patterns without necessitating additional trainable parameters.
 Figure 5: Instruction tuning loss and METEOR visualization.
 Figure 6: Model Scaling Performance on MIMIC-IV-ECG and PTB-XL.
-5.3.2 Scalability Analysis.
-To investigate whether ECG instruction tuning on larger-scale models yields better results, we validated LLaMA-2 models of 7B and 13B parameter sizes on both MIMIC-IV-ECG and PTB-XL datasets.
-As depicted in 7, an upward trend in all evaluation metrics is observed with a gradual increase in model size.
+Scalability Analysis.
+To investigate whether ECG instruction tuning on larger-scale models yields better results, we validated LLaMA-2 models of 7B, 13B, and 70B parameter sizes on both MIMIC-IV-ECG and PTB-XL datasets.
+As depicted in Figure 6, an upward trend in all evaluation metrics is observed with a gradual increase in model size.
 However, it is noteworthy that the gains in performance associated with increasing model size are not particularly significant.
@@
 Similarly, on the MIMIC-IV-ECG dataset, the 70B model’s F-1 score is only 0.01 higher than that of the 13B model.
-Therefore, we conjecture that enhancing both data scale and model size concurrently is necessary to achieve superior performance (Wei et al., 2022).
-5.3.3 Ablation Study on ECG Instruction Tuning.
-To validate the effectiveness of adopting an instruction tuning paradigm for aligning ECG signals and report representations, we conducted the following ablation experiment.
-We selected BLOOM, OPT, LLaMA-1, and Mistral as LLM backbones and removed the instruction tuning strategy, allowing the language models to fine-tune directly in the setting of generating text from ECG signals only.
-As shown in Figure 5, aligning ECG and report representations using direct fine-tuning, as opposed to instruction tuning, results in a significant decline across all metrics.
-Particularly in the Mistral model, the instruction tuning paradigm notably outperforms direct fine-tuning.
-These results demonstrate that, compared to fine-tuning, instruction tuning more effectively enhances the generalization ability of LLMs on unseen tasks/data (Ouyang et al., 2022).
-Figure 8: Examples of ECG reports generated by LLaMA-2 and Mistral-Instruct. We highlight the consistent information between the generated reports and the ground truths with blue color.
-5.3.4 Qualitative Results.
+Therefore, we conjecture that enhancing both data scale and model size concurrently is necessary to achieve superior performance [34].
+Figure 7: Ablation Study of ECG Instruction Tuning on MIMIC-IV-ECG Dataset.
+Ablation Study on ECG Instruction Tuning.
+We conducted an ablation study to evaluate instruction tuning’s impact on aligning ECG signals with report representations.
+Utilizing LLMs such as BLOOM, OPT, LLaMA-1, and Mistral without instruction tuning, we allowed direct learning from ECG signals.
+The findings, illustrated in Figure 7, indicate a significant performance drop across all metrics without instruction tuning, particularly in Mistral.
+This underscores instruction tuning’s superiority in enhancing LLMs’ generalization to new tasks/data over direct fine-tuning [35].
+Qualitative Results.
 In Figure 8, we randomly select two samples generated by MEIT using LLaMA-2 and Mistral-Instruct as the LLM backbones.
@@
 These generated reports demonstrate the efficacy of our method.
-6 Related Work
-Instruction Tuning.
-Instruction tuning (Zhang et al., 2023a; Wang et al., 2023) is designed to enhance the zero-shot capabilities of LLMs on new tasks using either machine-generated or human-annotated instructions.
-Notable examples such as InstructGPT (Ouyang et al., 2022), FLAN-PaLM (Chung et al., 2022), and Alpaca (Taori et al., 2023) utilize various methods, including manual collection and human feedback, to generate instruction template data for fine-tuning.
-In the realm of multimodal learning, models like LLaVA (Liu et al., 2023b), MiniGPT-4 (Zhu et al., 2023), and AnyMAL (Moon et al., 2023) demonstrate the utility of multimodal instruction data to enhance learning across different modalities.
-However, medical signals such as ECG and their corresponding text remain underexplored.
-Our work bridges this gap by developing a specialized ECG-related instruction tuning framework, thereby improving the generation of ECG reports.
+6 Related Works
+Instruction Tuning. Instruction tuning [36, 37] boosts zero-shot learning in LLMs for new tasks using instructions. Notable models like InstructGPT [35], FLAN-PaLM [38],
+Figure 8: Examples of ECG reports generated by LLaMA-2 and Mistral-Instruct. We highlight the consistent information between the generated reports and the ground truths with blue color.
+and Alpaca [39] fine-tune with instruction data through various methods, including human feedback.
+Similarly, multimodal models such as LLaVA [33], MiniGPT-4 [40], and AnyMAL [41] benefit from multimodal instructions for enhanced learning.
+However, instruction tuning for medical signals like ECG remains underexplored.
+Our research develops a targeted instruction-tuning framework for ECG report generation, addressing this gap.
 Medical Report Generation.
-Different from medical vision-language alignment (Wan et al., 2023a), medical Report Generation can be categorized into three main approaches: (1) Template Selection and Generation, where methods like HRGR (Li et al., 2018) focus on sentence creation or template retrieval, and CMAS (Jing et al., 2017) uses dual writers for differentiating regions; (2) Data Integration and Coherence, illustrated by PPKED’s (Liu et al., 2021b) knowledge merging for consistency, and CA’s contrasting of abnormal features (Ma et al., 2021); (3) Cross-Modal Alignment, with (Chen et al., 2022b; Qin & Song, 2022) using attention mechanisms for text-visual feature correlation.
-Recent advancements (Gu et al., 2024; Tanida et al., 2023; Li et al., 2023c) focus on local visual feature learning and multi-level cross-modal alignment, showcasing the evolution of semantic matching in radiology report generation.
-However, these approaches mainly focus on radiology report generation, directly applying these approaches to ECG data is challenging because the patterns between medical images and ECG biomedical signals are significantly different.
-Signal-to-Text Generation.
-Signal-to-Text (S2T) generation, predominantly speech-to-text, converts speech signals into text.
-Extensive research (Baevski et al., 2020; Chen et al., 2020; Liu et al., 2021a; Zheng et al., 2021; Hsu et al., 2021; Bapna et al., 2021) has extensively studied patterns in speech signals, analyzing both temporal and spectral aspects.
-Although the field of ECG report generation is largely unexplored, the use of speech signal research techniques in the creation of ECG reports presents significant challenges due to fundamental differences: (1) Speech signals for S2T tasks typically have a sampling frequency of up to 16,000 - 60,000 Hz (Lavry, 2012; Painter & Spanias, 2000; Byrne et al., 1994) and are monochannel, while ECG signals usually have a lower sampling frequency of around 500 Hz and encompass 12 channels.
-(2) Speech signals can be discretized into acoustic units corresponding to specific words (Défossez et al., 2022), but ECG signals lack such discrete units.
-Certain cardiac conditions, like bundle branch block, can obscure or distort the heartbeats patterns in ECG readings.
-(3) Speech signals and their transcriptions maintain strong temporal alignment, as each word correlates with a specific acoustic pattern, and the transcription follows the temporal sequence of speech (Zhang et al., 2023b).
-In contrast, ECG reports evaluate the cardiac patterns on a global level, assessing the entire ECG period rather than focusing on individual acoustic-like units in speech.
-These fundamental differences between speech-text pairs and ECG records, in both signal and text terms, highlight the need for S2T methods specifically designed for ECG data.
-Furthermore, the scarcity of research in the generation of ECG reports highlights an urgent need for comprehensive benchmarks using large-scale, publicly accessible data sets.
+Medical report generation [42]involves three strategies: (1) Template Selection and Generation, highlighted by HRGR [43] and CMAS [44]; (2) Data Integration and Coherence, as seen in PPKED [45] and CA [46]; (3) Cross-Modal Alignment, with efforts like [47, 48].
+While recent research has advanced these areas, applying them to ECG data is challenging due to the distinct nature of ECG signals compared to medical images.
+LLM for ECG Understanding.
+Only a few research efforts have focused on utilizing LLMs for ECG-related tasks [49, 50, 51].
+In the studies by [49, 51], the authors manually convert ECG signals into text features before feeding them into LLMs, rather than using the original ECG data.
+This approach overlooks the modality patterns inherent in the signals.
+Moreover, these studies only employ LLMs for classifying diseases from ECG data and do not explore the generation of medical reports.
+In contrast, [50] attempts ECG report generation by using handcrafted ECG features as inputs instead of the original signals.
+However, they have not publicly made any code or models available, making direct comparisons challenging.
 7 Conclusion and Future Work
@@
 In future work, we plan to extend our framework to encompass additional medical signal domains, such as EEG signals, with the aspiration that our work spurs further advancements in developing more proficient medical-signal LLMs.
-Broader Impact
-Our MEIT framework first attempt to addresses automatic ECG report generation through ECG instruction tuning, establishing the first comprehensive benchmark for this process with mainstream LLMs.
-In this work, we mainly focus on generating reports from multimodal ECG instructions using LLMs.
-However, the generated results are not fully explainable or controllable, even though the generation procedure is transparent and trackable.
-This is because the underlying theory of LLMs remains largely unexplored, necessitating further investigation to ensure the quality and safety of the generated content.
-In the future, we aim to enable LLMs to utilize external, expert-verified knowledge databases, such as clinical protocols and medical textbooks, to enhance the explainability of the generated ECG reports.
-Appendix A Hyper-parameters of ECG Instruction Tuning
+Appendix A Appendix.
+A.1 Limitations and Future Work
+In this work, we propose an effective and efficient method for generating ECG reports.
+Since this is a generative task, the content produced may occasionally lead to hallucinations.
+This is particularly pertinent in medical contexts such as ECG report generation, where accuracy is of utmost importance.
+To mitigate this issue, future iterations of our work could incorporate retrieval-augmented methods [14], information optimized strategies [52, 53, 54].
+For instance, by integrating a retrieval mechanism that accesses a database of verified medical or biological knowledge [55] or previous ECG reports, the model could be guided towards generating more accurate and reliable outputs.
+A.2 Hyper-parameters of ECG Instruction Tuning
 Table 5: Hyper-parameters of ECG instruction tuning for all LLM backbones.
@@
 Mistral-Instruct 128
-In this study, we concentrate on the implementation of the Low-Rank Adaptation (LoRA) (Hu et al., 2021) technique for efficient fine-tuning, specifically applied to ECG instruction tuning.
-As detailed in Table A provided, we utilize mixed precision at bf16 for enhanced computational efficiency.
+In this study, we implement the Low-Rank Adaptation (LoRA) [16] technique for efficient fine-tuning, specifically applied to ECG instruction tuning.
+As detailed in Table A.2 provided, we utilize mixed precision at bf16 for enhanced computational efficiency.
 Our models undergo instruction tuning over 5 epochs, with LoRA parameters set at an alpha of 64 and a rank of 128, accompanied by a dropout rate of 0.1.
-The total batch size is 64 with a gradient accumulation factor of 2.
+The total batch size is 64, with a gradient accumulation factor of 2.
 The maximum sequence length is constrained to 256 tokens.
-Additionally, we adopt a learning rate with 2e-5 for GPT-NeoX, and 1e-4 for the other models, optimized using the AdamW algorithm.
+Additionally, we adopt a learning rate with 2e-5 for GPT-NeoX and 1e-4 for the other models, optimized using the AdamW algorithm.
 The learning rate follows a linear schedule with a warm-up ratio of 0.03.
 We set the weight decay to 0.0.
-Moreover, as shown in Table A, we detail the ECG embedding dimensions for various language models, highlighting their approach to ECG data encoding.
-GPT2-Medium and GPT2-Large feature ECG dimensions of 64, while GPT-Neo, BLOOM, OPT, LLaMA-1, Mistral, LLaMA-2, and Mistral-Instruct use a dimension of 128.
+Moreover, as shown in Table A.2, we detail the ECG embedding dimensions for various language models, highlighting their approach to ECG data encoding.
+GPT2-Medium and GPT2-Large feature ECG dimensions 64, while GPT-Neo, BLOOM, OPT, LLaMA-1, Mistral, LLaMA-2, and Mistral-Instruct use a dimension of 128.
 GPT-NeoX employs a dimension of 96, and GPT-J notably uses the largest dimension of 256.
 These dimensions, reflecting each model’s head dimension design, illustrate diverse strategies in ECG data processing across different models.
-Appendix B Visualization of Generated ECG Report Samples
+A.3 More Details of ECG Encoder
+Projection Layer For the design of the projection layer within the ECG encoder, we adopt a non-linear approach similar to CLIP [56] and Med-UniC [57].
+Specifically, in our experiments, we employ two consecutive linear layers, each followed by BatchNorm1d [^4] https://pytorch.org/docs/stable/generated/torch.nn.BatchNorm1d.html.
+Besides, ReLU serves as the activation function between the two linear layers.
+The default settings for input and hidden layers dimensions are set to 2048 in our experiment.
+Parameter Size Analysis To demonstrate the ECG encoder’s lightweight design, we analyzed its trainable parameters during instruction tuning and total parameters during inference, using the LLaMA-1 7B model for illustration (Table 7).
+The analysis reveals the ECG encoder’s trainable parameters are substantially fewer than those of the LoRA adapter in the LLM backbone during instruction tuning, and its parameter share of the overall framework is minimal for inference, underscoring its efficiency.
+Table 7: Parameter Comparison of ECG encoder and LLM backbone. We use LLaMA-1 7B as an example.
+Module Trainable Params Inference Params
+LLM backbone 159M 6.90B
+ECG encoder 20.4M 20.4M
+A.4 Further Analysis of Generated Prompts
+In the ECG instruction data curation, we manually created 32 prompt examples, as illustrated in Section 3.1.
+To increase the diversity of our samples, we employed GPT-4 to rephrase these manually designed prompts, generating a larger pool of prompt examples.
+These generated examples were randomly sampled and paired with ECG-text pairs to compile the ECG instruction dataset.
+In this section, We compare the experiment’s effects using 128, 256, and 512-generated samples, respectively.
+Table 8 shows the corresponding results with different dimensions.
+When the number is 256, it can achieve better results in most experimental settings.
+Hence, we take 256 generated samples as our default setting during the instruction tuning and inference.
+Table 8: Performance comparison of different numbers of generated prompt samples.
+We evaluate them on the MIMIC-IV-ECG dataset, using BLEU-4, METEOR, ROUGE-L, and CIDEr-D metrics.
+We take LLaMA-1 7B as the LLM backbone here.
+heavy teal color indicates the highest results.
+Prompt Nums BLEU-4 METEOR ROUGE-L CIDEr-D
+128 0.541 0.756 0.718 5.15
+256 0.543 0.761 0.724 5.26
+512 0.538 0.754 0.732 5.03
+A.5 Visualization of Generated ECG Report Samples
 As illustrated in Figures 9,  10, and  11 we have visualized the report samples generated by LLaMA-1, LLaMA-2, and Mistral-Instruct.
 The samples are presented in blue font to highlight the key information that aligns with the ground truth.
-The visualization demonstrates that all three selected models are capable of capturing the essential patterns of ECG signals and generating accurate reports.
-This underscores the efficacy of our proposed MEIT framework, which proves to be adaptable to the majority of open-source LLMs.
+The visualization demonstrates that all three selected models can capture the essential patterns of ECG signals and generate accurate reports.
+This underscores the efficacy of our proposed MEIT framework, which is adaptable to most open-source LLMs.
 It effectively learns the correct clinical semantics of ECG signals, thereby enabling the generation of corresponding reports.

### v4

@@
-MEIT: Multi-Modal Electrocardiogram Instruction Tuning on Large Language Models for Report Generation
+MEIT: Multimodal Electrocardiogram Instruction Tuning on Large Language Models for Report Generation
 Abstract
@@
 Our approach uniquely aligns the representations of the ECG signal and the report, and we conduct extensive experiments to benchmark MEIT with nine open-source LLMs using more than 800,000 ECG reports.
-MEIT’s results underscore the superior performance of instruction-tuned LLMs, showcasing their proficiency in quality report generation, zero-shot capabilities, and resilience to signal perturbation.
-These findings emphasize the efficacy of our MEIT framework and its potential for real-world clinical application.
+MEIT’s results underscore the superior performance of instruction-tuned LLMs, showcasing their proficiency in quality report generation, zero-shot capabilities, resilience to signal perturbation, and alignment with human expert evaluation.
+These findings emphasize the efficacy of MEIT and its potential for real-world clinical application.
+Our code is released at https://github.com/AIoT-MLSys-Lab/MEIT.
+MEIT: Multimodal Electrocardiogram Instruction Tuning on
+Large Language Models for Report Generation
+Zhongwei Wan1, Che Liu2, Xin Wang1, Chaofan Tao3, Hui Shen1, Jing Xiong3, Rossella Arcucci2, Huaxiu Yao4, Mi Zhang1 1The Ohio State University  2 Imperial College London 3The University of Hong Kong  4University of North Carolina at Chapel Hill https://github.com/AIoT-MLSys-Lab/MEIT
 1 Introduction
-Diagnosing heart diseases is of paramount importance, requiring the analysis of multimodal data sources, which include electrocardiogram (ECG) recordings and clinical reports.
-Cardiologists read and interpret these ECG recordings to manually generate comprehensive ECG reports for heart disease diagnosis, which is a complex and time-consuming process.
-Recently, AI models have been developed to facilitate ECG data analysis for the task of classification [1, 2, 3, 4].
-Despite these efforts, the automatic generation of reports from ECG recordings still needs to be explored.
-Unlike other AI-empowered medical report generation applications (e.g., radiology reports), the primary challenge for ECG report generation stems from the distinct nature of ECG content.
-ECG reports, often comprising brief phrases that summarize signal patterns, contrast with detailed anatomical descriptions in radiology reports.
-The difference in the content and semantic interpretation between imaging and ECG data complicates the direct application of radiology-focused methods to ECG reports.
+Electrocardiogram (ECG) is the primary tool for heart disease diagnosis.
+In standard practice, cardiologists examine these ECG recordings and manually generate detailed diagnostic reports, which is a complex and time-consuming process.
+Recently, AI models have been developed to streamline ECG data analysis for the classification task (Hu et al., 2023; Liu et al., 2023a, 2024a), yet the automatic generation of reports from ECG recordings remains relatively underexplored.
+Unlike image-based medical report generation tasks (e.g., radiology reports), ECG report generation faces unique challenges due to the concise, keyword-centric content of ECG reports, which contrasts with the more extensive anatomical descriptions in radiology.
+Directly transferring methods from radiology is hindered by the distinct nature of ECG signals and the limited semantic overlap with imaging data.
 Furthermore, there is still a lack of comprehensive benchmarks for evaluating the performance of ECG report generation.
-To tackle these challenges, we introduce MEIT, a Multimodal ECG Instruction Tuning framework that extends the capabilities of LLMs in the cardiology context to generate ECG reports using ECG recordings and human instructions.
-Inspired by the versatility of LLMs [5, 6, 7, 8, 9] in handling diverse language tasks simultaneously, we develop a specialized instruction tuning process for ECG report generation.
-MEIT aligns human instructions with ECG recordings, enabling LLMs to generate professional-grade reports and exhibit zero-shot report generation capabilities under domain transfer scenarios across various continents and data collection devices.
-Specifically, leveraging publicly available ECG datasets, we construct a multimodal instruction dataset including ECG records, human instructions, and paired reports.
-Then, we propose an effective and efficient attention-based fusion method to integrate ECG and text representations in the latent space.
-This enables LLMs to understand ECG signals for report generation without introducing additional training parameters in the attention layer.
-In addition to the ECG report generation approach, we introduce a comprehensive benchmark for ECG report generation evaluation, utilizing two datasets with 20K and 800K ECG-report pairs, respectively, across three evaluation tasks: report generation quality, zero-shot learning across datasets, and robustness analysis in the face of ECG signal perturbation.
-Utilizing the ECG report evaluation benchmark, we assess the proposed approach across nine open-source LLMs.
-The results demonstrate (1) the superior performance of MEIT in ECG report generation and the effective learning and alignment of ECG representations; (2) the effective transferability of LLMs under the MEIT framework in domain transfer scenarios.
-To summarize, our primary contribution is the MELT framework, a novel approach to automating ECG report generation and evaluation based on LLMs.
-This framework incorporates a lightweight, attention-based fusion module across various LLM models.
-Furthermore, we design a new benchmark for ECG report generation, which contains three evaluation tasks.
-Our evaluations showcase the enhanced capabilities of instruction-tuned LLMs in generating ECG reports, highlighting the transferability in zero-shot tests and robustness against data perturbations.
-MEIT paves the way for future advances in automated ECG report generation and methodological innovations in integrating biomedical signals into LLMs.
-2 Preliminaries
+Drawing on the versatility of LLMs and multimodal LLMs (MLLMs) (Achiam et al., 2023; Touvron et al., 2023b; Wan et al., 2023b; Wang et al., 2024a; Zhu et al., 2024b; Wang et al., 2024b; Zhang et al., 2025a; Zhu et al., 2025) in various language tasks, in this work, we introduce MEIT, a Multimodal ECG Instruction Tuning framework that extends the capability of large language models (LLMs) for the task of ECG report generation.
+By aligning ECG recordings with human instructions, MEIT produces clinically relevant reports and demonstrates zero-shot performance when transferring across different continents and data collection devices.
+Concretely, we construct a multimodal instruction dataset from publicly available ECG recordings and propose an efficient attention-based fusion method that incorporates ECG signals into the latent space of LLMs without adding new parameters.
+We also introduce a comprehensive benchmark to evaluate ECG report generation across two large-scale datasets (with 20K and 800K ECG-report pairs), covering four tasks: (1) report generation quality, (2) zero-shot transfer across datasets, (3) robustness under signal perturbations, and (4) alignment with expert evaluation.
+We assess MEIT using ten open-source LLMs, demonstrating: (1) the superior performance of MEIT in ECG report generation and effective ECG representation learning; and (2) the strong transferability of instruction-tuned LLMs across diverse clinical domains.
+In summary, our primary contribution is the MEIT framework, a novel method for automated ECG report generation and evaluation based on LLMs.
+It features a lightweight, attention-based fusion module for integrating ECG signals and text, along with a newly designed four-task benchmark for ECG report evaluation.
+Empirical results highlight MEIT’s ability to generate high-quality reports, perform robustly under data perturbations, and align closely with expert assessments, thereby paving the way for improved ECG interpretation and broader innovations in embedding biomedical signals into LLMs.
+2 Related Work
+Medical Report Generation.
+Our work is closely related to medical report generation, which has been extensively studied in the context of medical images.
+Existing methods can be grouped into three main categories: (1) template-based methods, such as HRGR (Li et al., 2018) and CMAS (Jing et al., 2017); (2) data integration and coherence methods, such as PPKED (Liu et al., 2021) and CA (Ma et al., 2021); and (3) cross-modal alignment methods, exemplified by Chen et al.
+(2022) and Qin and Song (2022).
+However, these methods are designed for medical images and struggle with ECG signals due to their distinct temporal and waveform characteristics.
+Instruction Tuning.
+Our work is also related to instruction tuning, which improves zero-shot learning in LLMs through task-specific instructions (Zhang et al., 2023; Wang et al., 2023).
+Models like InstructGPT (Ouyang et al., 2022), FLAN-PaLM (Chung et al., 2022), and Alpaca (Taori et al., 2023) leverage instruction data, including human feedback.
+Multimodal variants, such as LLaVA (Liu et al., 2023c), MiniGPT-4 (Zhu et al., 2023), and AnyMAL (Moon et al., 2023), extend this to visual tasks.
+However, these methods focus on natural images and are not directly applicable to ECG signals.
+To bridge this gap, we introduce a specialized instruction-tuning framework for ECG report generation.
+LLMs for ECG.
+Few studies have explored the use of LLM for the analysis of ECG signals (Liu et al., 2023b; Qiu et al., 2023; Yu et al., 2023).
+Some of them (Liu et al., 2023b; Yu et al., 2023) convert ECG signals into text features before importing them into LLMs, bypassing crucial modality-specific characteristics.
+In addition, these works focus on the task of disease classification rather than report generation.
+In contrast, our method directly processes ECG signals and focuses on the task of report generation.
+Figure 1: (a) Overview of MEIT; (b) Illustration of the architecture of Report Generator.
+3 MEIT
+3.1 Preliminaries
 Electrocardiogram (ECG) measures the electrical activity of an individual’s heart over time.
-An ECG recording typically contains a 12-lead multivariate time series, which acts as a 12-dimensional sequence of embeddings.
-The ECG signal offers a comprehensive view, encompassing both spatial and temporal aspects of cardiac function.
-ECG leads can be categorized into six limb leads (i.e., I, II, III, aVR, aVL, and aVF) to monitor arms and legs, providing frontal plane views, and six precordial leads (i.e., V1, V2, V3, V4, V5, and V6) to monitor chest, showing horizontal plane views.
-We denote an ECG recording as $\mathbf{X}_{e}\in\mathcal{R}^{M\times T}$, where $M$ represents the number of leads, and $T$ is signal length.
-Each ECG recording is associated with an ECG report $\mathbf{X}_{t}$ for description.
-Thus, we denote each ECG pair as $\{\mathbf{X}_{e},\mathbf{X}_{t}\}$.
-More details on visualization can be found in the appendix A.5.
-Figure 1: Illustration of the proposed MEIT framework, including ECG instruction data curation, instruction tuning on auto-regressive objective, and inference to get the generated report.
-Figure 2: Illustration of model architecture for ECG Report Generation.
-$\mathbf{K}_{e}$ and $\mathbf{V}_{e}$ refer to linear projection of $\mathbf{H}_{e}$ by multiplying shared $\mathbf{W}_{k}$ and $\mathbf{W}_{v}$ in the attention layer.
-3 MEIT
-Figure 2 illustrates the proposed MEIT framework.
-We begin by describing the curation process for ECG instruction data.
-We then delve into the architecture of the multimodal LLM designed for ECG report generation.
-Lastly, we describe the approach to generating ECG reports with instruction tuning.
-3.1 ECG Instruction Data Curation
-Given an ECG signal $\mathbf{X}_{e}$, our goal during inference is to generate an ECG report using an instruction prompt.
-For instance, the prompt can be “Given the ECG signal embeddings, please help me generate an accurate description for this ECG signal embeddings: ".
-To achieve this goal, during the training phase, we aim to create instruction tuning data to generate a response $\mathbf{\hat{X}}_{t}$ that aligns semantically with the ground truth $\mathbf{X}_{t}$.
-In addition, since we cannot predict the exact instruction prompt that users will use, we need to ensure that our report generation process is robust enough to handle different prompts.
-To address this challenge, we manually design some prompt samples, then utilize GPT-4 [5] to generate a set of prompts by rephrasing, as shown in Figure 2.
-Then, we randomly select one instruction prompt $\mathbf{X}_{p}$ from the prompt set and create a general instruction-following template: <|user|>: $\mathbf{X}_{p}$, $\mathbf{X}_{e}$}<|assistant|>: $\mathbf{X}_{t}$} </s>, where <|user|> and <|assistant|> are added special tokens for tokenizer, </s> is a stop sign for each response.
+An 12-lead ECG recording is a multivariate time series.
+It offers a multi-dimensional view, encompassing both spatial and temporal aspects of cardiac function.
+The 12 ECG leads include six limb leads (i.e., I, II, III, aVR, aVL, and aVF) that monitor arms and legs, providing frontal plane views, and six precordial leads (i.e., V1, V2, V3, V4, V5, and V6) that monitor chest, showing horizontal plane views.
+Formally, let $\mathbf{X}_{e}\in\mathcal{R}^{M\times T}$ denote an ECG recording, where $M$ represents the number of leads, and $T$ represents the signal length.
+Each ECG recording is associated with an ECG report $\mathbf{X}_{t}$ that describes and interprets the ECG recording.
+Thus, we denote each ECG recording-report pair as $\{\mathbf{X}_{e},\mathbf{X}_{t}\}$.
+3.2 Overview of MEIT
+Figure 1 (a) illustrates the proposed MEIT framework.
+In the data curation stage, we construct the ECG instruction tuning data, which includes instruction prompts, ECG recordings and the corresponding ground truth ECG reports.
+In the instruction tuning stage, the ECG instruction tuning data are fed into the Report Generator for training using an autoregressive approach.
+During inference, the instruction prompts and the ECG recordings are inputted into the Report Generator to generate the ECG reports.
+In the following sections, we describe each component in detail.
+3.3 Data Curation
+Given an ECG recording $\mathbf{X}_{e}$, our goal during inference is to generate an ECG report using an instruction prompt such as “Given the ECG recording, please help me generate a report for this ECG recording:".
+To achieve this goal, we aim to create a set of instruction tuning data to generate a response $\mathbf{\hat{X}}_{t}$ that semantically aligns with the ground truth $\mathbf{X}_{t}$.
+Since we cannot predict the exact instruction prompt that users will use, we need to ensure that our report generation process is robust enough to handle different prompts.
+To do so, we manually design a small set of prompt samples, and then utilize GPT-4 (Achiam et al., 2023) to generate a larger prompt set by rephrasing, as shown in Figure 1 (a).
+For each ECG recording-report pair as $\{\mathbf{X}_{e},\mathbf{X}_{t}\}$, we randomly select one instruction prompt $\mathbf{X}_{p}$ from the larger prompt set and create an instruction tuning template: <|user|>: $\mathbf{X}_{p}$, $\mathbf{X}_{e}$} <|assistant|>: $\mathbf{X}_{t}$} </s>, where <|user|> and <|assistant|> are added special tokens for tokenizer, and </s> is a stop sign for each response.
 This approach ensures that the generated response conveys the same meaning as the ground truth and remains adaptable to different instruction prompts.
-Following this strategy, we construct the ECG instruction data using a MIMIC-IV-ECG [10] dataset that contains 80K annotated data and a 20K dataset PTB-XL [11].
-The ECG instruction data samples are shown in Appendix A.5.
-3.2 ECG Report Generation Model
-In MEIT, the multimodal ECG report generation model decodes ECG signals end-to-end to generate ECG reports.
-The architecture is illustrated in Figure 2.
-Specifically, the report generation model directly encodes an entire ECG-signal $\mathbf{X}_{e}\in\mathcal{R}^{M\times T}$ into latent embeddings and integrate it with the language embeddings with modality alignment, and then autoregressively generate the ECG report.
-Next, we detail each component of the report generation model.
+Following this strategy, we construct the ECG instruction tuning data using the PTB-XL (Wagner et al., 2020) dataset and the MIMIC-IV-ECG (Gow et al., ) dataset.
+Some examples of ECG instruction tuning data are included in Appendix A.8.
+3.4 Report Generation
+Figure 1 (b) illustrates the architecture of the Report Generator.
+As shown, the Report Generator utilizes an ECG encoder to encode $\mathbf{X}_{e}$ into ECG embeddings and integrates them with the language embeddings extracted from the corresponding instruction tuning template through an ECG-text alignment module to autoregressively generate the ECG report.
+In this section, we describe the key components of the Report Generator in detail.
 ECG Encoder.
-Since the ECG signal is of high resolution in the temporal domain, it is vital to efficiently extract temporal features per lead before interaction with semantic embeddings inside the LLM backbone.
-Our default ECG encoder $\mathcal{F}_{e}(\cdot)$ consists of temporal convolution blocks to encode each ECG signal into embeddings.
+Since the ECG recording is high resolution in the time domain, it is vital to efficiently extract temporal features per lead before interaction with semantic embeddings inside the LLM backbone.
+Our default ECG encoder $\mathcal{F}_{e}(\cdot)$ consists of temporal convolution blocks to encode the ECG recordings into embeddings.
 Specifically, each temporal convolution block comprises several 1-D convolution layers, batch normalization layers, and ReLU activation layers, followed by average pooling.
-To align the output dimension with the head dimension of LLM backbone $\mathcal{F}_{l}(\cdot)$, we employ a non-linear projection layer $\mathcal{P}_{e}(\cdot)$ to obtain the ECG embeddings:
+This design allows us to effectively capture temporal dependencies and reduce the complexity of the signal representations, ensuring that the model can quickly learn important temporal features efficiently.
+To further align the output dimension with the head dimension of the LLM backbone $\mathcal{F}_{l}(\cdot)$, we employ a non-linear projection layer $\mathcal{P}_{e}(\cdot)$ to generate the ECG embeddings:
 $\mathbf{H}_{e}=\mathcal{P}_{e}\left(\mathcal{F}_{e}\left(\mathbf{X}_{e}\right)\right),$ (1)
 where $\mathbf{H}_{e}\in\mathcal{R}^{D_{h}}$, $D_{h}$ has the same dimension as the multi-head attention layers of LLMs.
-Note that our default ECG encoder is lightweight and is able to learn temporal patterns of signals without a long training period.
-More details about ECG Encoder are illustrated in Appendix A.3.
-ECG Modality Alignment.
-We introduce an ECG modality alignment strategy to guide the LLMs in aligning ECG signal data with corresponding textual outputs.
-This approach is detailed in Figure 2.
-Specifically, given the ECG embeddings $\mathbf{H}_{e}$, the alignment strategy incorporates $\mathbf{H}_{e}$ with the current hidden state $\mathbf{H}_{t}^{i}$ generated from previous $i-1$-th layer of the LLM backbone $\mathcal{F}_{l}\left(\cdot\right)$ for next-token prediction task.
+Note that our default ECG encoder is lightweight and is able to learn temporal patterns of ECG recordings without a long training period.
+More details about the ECG encoder are included in Appendix A.3.
+ECG-Text Alignment.
+Given the ECG and the text embeddings, we introduce an ECG-text alignment strategy to guide the LLM in aligning the ECG embeddings with the text embeddings.
+As shown in Figure 1 (b), given the ECG embeddings $\mathbf{H}_{e}$, the ECG-text alignment strategy incorporates $\mathbf{H}_{e}$ with the current hidden state $\mathbf{H}_{t}^{i}$ generated from the previous ($i-1$)th layer of the LLM backbone $\mathcal{F}_{l}\left(\cdot\right)$ for next-token prediction task.
 Here $\mathbf{H}_{t}^{i}$ is defined as:
@@
 where $i$ is the current layer index.
-Different from gated-attention fusion in Flamingo [12], Memorizing Transformer [13] , and G-MAP [14], or Q-former in BLIP-2 [15] that require the additional trainable parameters and design complex modules for two-stage training, we propose a simple yet effective concatenated-fusion method that directly injects the ECG embeddings with language context in the causal-attention (C-Attn) of the models.
+Traditional gated-attention fusion methods like Flamingo (Alayrac et al., 2022), Memorizing Transformer (Wu et al., 2022), G-MAP (Wan et al., 2022), and Q-former in BLIP-2 (Li et al., 2023) require additional trainable parameters and are designed for complex multi-stage alignment of rich semantic information (e.g., images).
+Different from them, our method provides a lightweight concatenation-based alignment strategy tailored to the ECG embeddings, enabling efficient learning of ECG semantic features via directly injecting the ECG embeddings with language context in the self-attention, while preventing potential catastrophic forgetting of general knowledge in LLMs.
 In our approach, each attention layer combines $\mathbf{H}_{e}$, generated from the ECG encoder and projector as a prefix condition, with $\mathbf{H}_{t}^{i}$, derived from the preceding layer.
 The fusion process is as follows:
-$\operatorname{C-Attn}\left(\mathbf{H}_{e},\mathbf{H}_{t}^{i}\right)=\left[\operatorname{head}_{1},\ldots,\operatorname{head}_{k}\right]\mathbf{W}_{o},$ (3)
+$\operatorname{Self-Attn}\left(\mathbf{H}_{e},\mathbf{H}_{t}^{i}\right)=\left[\operatorname{head}_{1},\ldots,\operatorname{head}_{k}\right]\mathbf{W}_{o},$ (3)
 where $k$ represents the number of attention heads, and $\mathbf{W}_{o}$, a matrix in $\mathcal{R}^{kD_{h}\times D_{m}}$, serves as the projection matrix with $D_{m}$ denoting the hidden size of the LLM backbone.
@@
 $\mathbf{W}_{q,j}$, $\mathbf{W}_{k,j}$, and $\mathbf{W}_{v,j}$ in $\mathcal{R}^{D_{h}\times D_{h}}$ represent the projection matrices for query, key, and value for each head $j$, respectively.
-Our model’s design allows for the efficient fusion of two modalities through causal attention, facilitating conditional generation without the need for additional parameter updates to align the ECG modality.
-More comparisons about ECG modality alignment and other fusion approaches are illustrated in Appendix LABEL:appendix:_ECG_align_ablation.
-3.3 ECG Instruction Tuning
-As described in Section 3.1, we have converted ECG-text pairs into a chat-bot style instruction format: <|user|>: $\mathbf{X}_{p}$, $\mathbf{X}_{e}$}<|assistant|>: $\mathbf{X}_{t}$}</s>.
-During instruction tuning, we compute autoregressive loss only on tokens after response tokens <assistant>, and use label loss masking to finetune the model, where we mask all tokens belonging to $\mathbf{X}_{p}$ and $\mathbf{X}_{e}$.
-To save computational resources and accelerate the convergence of instruction tuning, we use LoRA [16] adapters for all linear layers of the LLM backbone $\mathcal{F}_{l}$ and freeze its backbone.
+Such design allows for the efficient fusion of two modalities through causal attention, facilitating conditional generation without the need for additional parameter updates to align the ECG modality with the text modality.
+Ablation studies comparing with other fusion methods demonstrate the effectiveness and efficiency of our proposed lightweight alignment strategy.
+More comparisons about ECG-text alignment and other fusion approaches are included in Table 5.
+3.5 Instruction Tuning
+During instruction tuning, we follow the instruction tuning template: <|user|>: $\mathbf{X}_{p}$, $\mathbf{X}_{e}$} <|assistant|>: $\mathbf{X}_{t}$} </s> and compute the autoregressive loss only on tokens after response tokens <assistant>, and use label loss masking to finetune the model, where we mask all tokens belonging to $\mathbf{X}_{p}$ and $\mathbf{X}_{e}$.
+To save computational resources and accelerate the convergence of instruction tuning, we use LoRA (Hu et al., 2021) adapters for all linear layers of the LLM backbone $\mathcal{F}_{l}$ and freeze its backbone.
 Subsequently, given a sequence of ECG instruction data, we compute the probability of the target response $\mathbf{X_{t}}$ as an autoregressive function:
 $p\left(\mathbf{X}_{t}\mid\mathbf{X}_{p},\mathbf{X}_{e}\right)=\prod_{i=j}^{L}p_{\boldsymbol{\theta}}\left(\mathbf{x}_{t,i}\mid\mathbf{X}_{p},\mathbf{X}_{e},\mathbf{X}_{t,<i}\right),$ (6)
-where $j$ is the start index after <assistant>, $\theta$ is the trainable parameters of LoRA and ECG encoder $\mathcal{F}_{e}$, $\mathbf{X}_{t,<i}$ is the response tokens before the current generation $\mathbf{x}_{t,i}$.
-4 Benchmarking ECG Report Generation
-In this section, we create a benchmark to evaluate the ECG report generation approach.
-Our benchmark includes two datasets (PTB-XL and MIMIC-IV-ECG) along with three ECG-related tasks, including (1) the quality of ECG report generation, (2) the zero-shot ability under domain transfer, (3) the robustness of the model against perturbed signals.
+where $j$ is the start index after <assistant>, $\theta$ is the trainable parameters of LoRA and ECG encoder $\mathcal{F}_{e}$, and $\mathbf{X}_{t,<i}$ is the response tokens before the current generation $\mathbf{x}_{t,i}$.
+4 ECG Report Generation Benchmark
 4.1 Datasets
+MIMIC-IV-ECG.
+MIMIC-IV-ECG (Gow et al., ) is the largest publicly available 12-lead ECG dataset.
+MIMIC-IV-ECG contains $800,035$ samples from $161,352$ subjects.
+Similar to PTB-XL, each sample includes a raw ECG recording and its report, sampled at $500$Hz for 10 seconds.
+The dataset is split into training, validation, and test sets at a ratio of 80%:10%:10% respectively.
 PTB-XL.
-The PTB-XL dataset [11] comprises 21,837 clinical 12-lead ECGs, each lasting 10 seconds, from 18,885 patients.
-Each ECG is accompanied by a corresponding report.
-The data includes raw ECG signals and their reports, with all recordings sampled at 500Hz over 10 seconds.
-We divided this dataset into training, validation, and testing subsets in a 70%:10%:20% ratio, respectively.
-The human experts double-check all samples in the test set to ensure data quality.
-As mentioned in Sec 3.1, we reformulate the dataset into the instruction data format.
-MIMIC-IV-ECG.
-The MIMIC-IV-ECG dataset [10] is currently the largest publicly accessible ECG dataset, containing 800,035 paired samples from 161,352 unique subjects.
-Similar to PTB-XL, each sample in this dataset includes a raw ECG signal and its corresponding report, with all recordings sampled at 500Hz for 10 seconds.
-The division of this dataset into training, validation, and testing subsets is 80%:10%:10% ratio.
-Likewise, we reconstruct this dataset into an ECG instruction data template.
+We also use the PTB-XL dataset (Wagner et al., 2020), which contains $21,837$ clinical 12-lead ECG recordings collected from $18,885$ patients.
+Each ECG recording is $10$ seconds long sampled at $500$Hz and has a corresponding report.
+We divide PTB-XL into training, validation, and test sets at a ratio of 70%:10%:20%, respectively.
+4.2 Models
+We use two smaller language models and ten LLMs based on the PEFT [^1] https://github.com/huggingface/peft library, which directly supports LoRA (Hu et al., 2021) to construct the multimodal ECG report generation model described in Section 3.4.
+The two smaller language models are GPT2-Medium and GPT-Large (Radford et al., 2019).
+The ten LLMs are GPT-Neo (Black et al., 2021), GPT-NeoX (Black et al., 2022), GPT-J (Wang and Komatsuzaki, 2021), BLOOM (Workshop et al., 2022), OPT (Zhang et al., 2022), LLaMA-1 (Touvron et al., 2023a), LLaMA-2-Instruct (Touvron et al., 2023b), LLaMA-3-Instruct (Touvron et al., 2023b), Mistral (Jiang et al., 2023), and Mistral-Instruct.
 4.3 Evaluation Metrics
-We evaluate models using nine metrics: BLEU 1-4 [17], METEOR [18], ROUGE 1-2 and L [19], CIDEr-D [20], and BERTScore [21].
-BLEU and METEOR assess machine translation quality, focusing on accuracy and fluency.
+We evaluate the performance using ten metrics: BLEU 1-4 (Papineni et al., 2002), METEOR (Banerjee and Lavie, 2005), ROUGE 1, 2 and L (Lin, 2004), CIDEr-D (Vedantam et al., 2015), and BERTScore (Zhang et al., 2019).
+Specifically, BLEU 1-4 and METEOR assess machine translation quality, focusing on accuracy and fluency.
 ROUGE-L measures sentence fluency and structure, while ROUGE-1 and ROUGE-2 examine uni-gram and bi-gram overlaps.
-CIDEr-D evaluates the relevance and uniqueness of generated ECG reports against a candidate set, and BERTScore assesses semantic similarity to the ground truth, ensuring content accuracy.
-4.3 Tasks
-We evaluate the performance of ECG report generation on the following three tasks:
-Quality Report Generation.
-This task aims to assess report quality after ECG instruction tuning using 10% of MIMIC-IV-ECG and PTB-XL datasets as the test set.
-The evaluation examines how closely generated reports match the original’s structure and meaning, considering various instructions and ECG inputs.
-We analyze metrics like BLEU-1 to 4, METEOR, ROUGE 1, 2, L, CIDEr-D, and BERTScore.
-Zero-shot Generalizability.
-To explore the generalizability of LLMs in domain transfer scenarios following ECG instruction tuning, we trained the models on 70% of the instruction data from MIMIC-IV-ECG.
-Following this, we evaluated the models’ zero-shot capabilities on the PTB-XL test set.
-It’s important to note that the PTB-XL and MIMIC-IV-ECG datasets originate from different continents—Europe and the United States, respectively—utilizing varied devices and from distinct hospitals, across different time periods.
-Therefore, we consider these datasets to represent two separate domains.
-This distinction allows us to use the PTB-XL dataset to gauge our model’s performance in zero-shot domain transfer effectively.
-We used the metrics BLEU-4, METEOR, ROUGE-L, and CIDEr-D because of limited space and calculated their average for model evaluation.
-Signal Perturbation Robustness.
-In real-world clinical settings, ECG signals often contain some degree of noise.
-To evaluate the robustness of MEIT against such noisy interference, we selected 10% of the ECG samples from the MIMIC-IV-ECG test dataset.
-We then added Gaussian noise to these samples during the models’ instruction-based inference process.
-For this evaluation, we used BLEU-4, METEOR, ROUGE-L, and CIDEr-D as metrics.
-5 Experiments and Evaluation
-5.1 Experimental Settings
-In this section, we evaluate and benchmark nine open-source decoder-only LLMs using the constructed ECG report generation benchmark.
-Additionally, we offer a comprehensive analysis of scalability and instruction tuning and present case studies showcasing the generated reports.
-Models.
-We use nine LLMs based on the peft [^1] https://github.com/huggingface/peft library, which directly supports LoRA [16] to construct the multimodal ECG report generation model described in Section 3.2.
-These models include GPT-Neo [22], GPT-NeoX [23], GPT-J [24], BLOOM [25], OPT [26], LLaMA-1 [27], LLaMA-2 [6], Mistral [28], and Mistral-Instruct [^2] https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.1, along with
+CIDEr-D evaluates the relevance and uniqueness of the generated ECG reports against a candidate set.
+Lastly, BERTScore assesses semantic similarity to the ground truth.
 Table 1: Natural language generation metric on MIMIC-IV-ECG.
 For model size, ’M’ denotes the million level, and ’B’ denotes the billion level.
-† refers to an improved instruct fine-tuned version of models.
+All checkpoints are downloaded from Hugging Face website.
+And all models have been fine-tuned using ECG instructions.
 The light teal color indicates the second highest results, and heavy teal color indicates the highest results.
@@
 Mistral 7B 0.697 0.659 0.611 0.571 0.763 0.740 0.765 0.658 5.48
-LLaMA-2† 7B 0.706 0.662 0.622 0.581 0.775 0.745 0.768 0.664 5.55
-Mistral-Instruct† 7B 0.714 0.665 0.619 0.576 0.768 0.751 0.762 0.667 5.62
+LLaMA-2-Instruct 7B 0.706 0.662 0.622 0.581 0.775 0.745 0.768 0.664 5.55
+Mistral-Instruct 7B 0.714 0.665 0.619 0.576 0.768 0.751 0.762 0.667 5.62
+LLaMA-3-Instruct 8B 0.733 0.686 0.648 0.610 0.799 0.773 0.795 0.686 5.78
 Table 2: Natural language generation metric on PTB-XL. The light teal color indicates the second highest results, and heavy teal color indicates the highest results.
@@
 Mistral 7B 0.486 0.475 0.446 0.421 0.673 0.591 0.697 0.634 3.98
-LLaMA-2† 7B 0.515 0.484 0.469 0.439 0.675 0.594 0.698 0.624 4.05
-Mistral-Instruct† 7B 0.501 0.481 0.457 0.425 0.664 0.592 0.700 0.641 4.01
-two relatively small pre-trained language models (GPT2-Medium and GPT-Large [29]) as fundamental baselines.
-Implementation Details. Our study utilized PyTorch 2.1, transformers [30], and accelerated on A100
+LLaMA-2-Instruct 7B 0.515 0.484 0.469 0.439 0.675 0.594 0.698 0.624 4.05
+Mistral-Instruct 7B 0.501 0.481 0.457 0.425 0.664 0.592 0.700 0.641 4.01
+LLaMA-3-Instruct 8B 0.539 0.513 0.494 0.467 0.698 0.615 0.725 0.646 4.45
+4.4 Tasks
+Task #1: Quality of the Generated Reports.
+In our first task, we aim to assess the quality of the generated ECG reports using 10% of PTB-XL and MIMIC-IV-ECG datasets as the test set.
+Specifically, this task examines how closely the generated reports match the ground truth reports in terms of structure and semantic meaning under various instructions and ECG recordings.
+Task #2: Zero-shot Generalizability.
+As the second task, to explore the generalizability of LLMs in domain transfer scenarios following ECG instruction tuning, we trained the models on 70% of the instruction data from MIMIC-IV-ECG.
+Following this, we evaluated the models’ zero-shot capabilities on the PTB-XL test set.
+It is important to note that the PTB-XL and MIMIC-IV-ECG datasets originate from different continents, Europe, and the United States, respectively, utilizing different devices and from different hospitals, across different time periods.
+Therefore, we consider these datasets to represent two separate domains.
+This distinction allows us to use the PTB-XL dataset to gauge our model’s performance in zero-shot domain transfer effectively.
+We used the metrics BLEU-4, METEOR, ROUGE-L, and CIDEr-D because of limited space and calculated their average for model evaluation.
+Task #3: Signal Perturbation Robustness.
+In real-world clinical settings, ECG signals often contain some degree of noise.
+To evaluate the robustness of MEIT against such noisy interference, we selected 10% of the ECG samples from the MIMIC-IV-ECG test dataset.
+We then added Gaussian noise to these samples during the models’ instruction-based inference process.
+For this evaluation, we used BLEU-4, METEOR, ROUGE-L, and CIDEr-D as metrics.
+Task #4: Evaluation of Alignment with Human Expert Annotations.
+Lastly, to evaluate the differences between the reports generated by MEIT and human expert annotations, we established specific evaluation criteria and utilized closed-source LLMs to conduct a professional assessment of both the generated reports and expert annotations.
+5 Experiments and Analysis
 Table 3: Semantic similarity between the generated ECG reports and ground truths is measured using BERTScore, denoted as P for Precision, R for Recall, and F-1 for the F-1 Score.
@@
 Mistral 0.761 0.732 0.746 0.711 0.664 0.687
-LLaMA-2† 0.764 0.725 0.744 0.721 0.668 0.693
-Mistral-Instruct† 0.773 0.722 0.747 0.730 0.661 0.694
-GPUs with LLMs from Hugging Face [31] ranging from 2.7 to 70 billion parameters.
+LLaMA-2-Instruct 0.764 0.725 0.744 0.721 0.668 0.693
+Mistral-Instruct 0.773 0.722 0.747 0.730 0.661 0.694
+LLaMA-3-Instruct 0.798 0.745 0.771 0.745 0.682 0.712
+5.1 Experimental Setup
+We utilized PyTorch 2.1, transformers (Wolf et al., 2020), and accelerated on A100 GPUs with LLMs from Hugging Face (Wolf et al., 2019) ranging from 2.7 to 70 billion parameters.
 For larger models, we used DeepSpeed [^2] https://github.com/microsoft/DeepSpeed.
@@
 Subsequently, we discard any samples whose reports contain fewer than 5 tokens.
-For, ECG encoder, we adopt random initialization.
-Additionally, the default number of generated prompts from GPT-4 is 256, more training, visualization details about ECG instruction tuning are illustrated in Appendix A.2, Section  5.3, and Appendix A.4.
-5.2 Benchmark Task Evaluation
-5.2.1 Quality Evaluation
+For ECG encoder, we adopt random initialization.
+Additionally, the default number of generated prompts from GPT-4 is 256.
+5.2 Quality Evaluation
 Performance on MIMIC-IV-ECG.
-Table 3 and 1 present the results of various types of language encoders $\mathcal{F}_{l}(\cdot)$ on MIMIC-IV-ECG.
-The results show that all LLMs perform better than smaller language models (SLMs), such as GPT2-Medium and GPT2-Large, across all evaluation metrics.
+Table 1 and 3 present the results of various types of language encoders $\mathcal{F}_{l}(\cdot)$ on MIMIC-IV-ECG.
+The results show that all ten LLMs perform better than the two smaller language models (GPT2-Medium and GPT2-Large) across all evaluation metrics.
 Notably, from GPT-Neo to Mistral-Instruct, LLM-based backbones achieve a significant margin over SLMs in all metrics.
 For instance, compared to GPT2-Large, the METEOR score increases in the range of 0.132 to 0.18 from GPT-Neo to LLaMA-2, and Mistral-Instruct outperforms GPT2-Large with an improvement of 0.18 in the ROUGE-L score and 0.134 in the F-1 of BERTScore.
-Figure 3: Zero-shot performance on PTB-XL dataset. “IT” denotes instruction tuning.
-The observed performance underscores the adeptness of LLMs in generalizing from signal data, showcasing enhanced proficiency in aligning ECG signal representations with corresponding textual information.
-This highlights the significant potential of LLMs in medical signal-to-text generation.
-Particularly, LLaMA-2 and Mistral-Instruct surpass their counterparts in most evaluative metrics, suggesting that models pre-tuned with general instructions are more adept at learning ECG-text alignment.
+The observed performance underscores the adeptness of LLMs in generalizing from ECG data showcasing enhanced proficiency in aligning ECG data representations with corresponding textual information.
+This highlights the significant potential of LLMs in ECG-to-text generation.
+Particularly, LLaMA-2-Instruct, Mistral-Instruct, and LLaMA-3-Instruct surpass their counterparts in most evaluative metrics, suggesting that models pre-tuned with general instructions are more adept at learning ECG-text alignment.
 Performance on PTB-XL.
@@
 Specifically, LLaMA-2 surpasses GPT2-Large by 0.134 in the BLEU-3 metric, while LLaMA-1 achieves a 0.103 improvement in the METEOR score.
-The overall experimental results reveal that Mistral-Instruct and LLaMA-2 are consistently the top two performers across most metrics because of their strong general instruction-following capabilities.
-5.2.2 Zero-shot Evaluation in Domain Transfer.
+The overall experimental results also reveal that Mistral-Instruct, LLaMA-2-Instruct, and LLaMA-3-Instruct are consistently the top two performers across most metrics because of their strong general instruction-following capabilities.
+Table 4: Evaluation results of LLaMA-2-Instruct and LLaMA-3-Instruct against human expert-annotated ground-truth reports. Each dimension is scored on a scale of 1 to 5.
+Model Medical Terminology Accuracy Logical Consistency Completeness Diagnostic Accuracy
+LLaMA-2-Instruct 4.25 4.11 3.72 3.60
+LLaMA-3-Instruct 4.52 4.38 4.01 3.98
+Figure 2: Zero-shot performance on PTB-XL dataset. “IT” denotes instruction tuning.
+5.3 Zero-shot Evaluation in Domain Transfer
+Although both PTB-XL and MIMIC-IV-ECG datasets are time-series data, they differ significantly in several aspects, including population (European vs.
+American), diverse collection devices, continents (Europe vs.
+US), protocols, and hospitals.
+These differences introduce substantial medical domain gaps (Bilheimer and Klein, 2010; Ross et al., 2020).
 In Figure 2, we present the evaluation of the zero-shot learning capabilities of various LLMs, which is trained on the MIMIC-IV-ECG dataset and then tested on PTB-XL (unseen dataset).
@@
 The involvement of ECG instruction tuning on MIMIC-IV-ECG enables the models to achieve superior zero-shot performance on the unseen PTB-XL dataset, indicating the necessity of instruction tuning in enhancing the models’ zero-shot ability on unseen datasets in ECG report generation.
-5.2.3 Robust Analysis with Perturbed ECG Signal.
-In a noise stress evaluation [32], we added Gaussian noise to ECG signals at signal-to-noise ratios (SNRs) of 0.05, 0.1, 0.15, and 0.2 during testing to assess model robustness.
-Our experiments utilized four LLM architectures: BLOOM, OPT, LLaMA-1, and Mistral, each trained on clean ECG signals from the MIMIC-IV-ECG
-Figure 4: Signal perturbation robustness analysis on various LLMs
-training set and tested on corresponding noise-added signals from its test set.
+5.4 Robust Analysis with Perturbed ECG Recordings
+Figure 3: Signal perturbation robustness analysis on various LLMs.
+In a noise stress evaluation (Wang et al., 2019), we added Gaussian noise to ECG signals at signal-to-noise ratios (SNRs) of 0.05, 0.1, 0.15, and 0.2 during testing to assess model robustness.
+Our experiments utilized four LLM architectures: BLOOM, OPT, LLaMA-1, and Mistral, each trained on clean ECG signals from the MIMIC-IV-ECG training set and tested on corresponding noise-added signals from its test set.
 The results, illustrated in Figure 3, show a performance decline in all LLMs as SNR decreased, highlighting the significant interference of ECG noise.
@@
 Even with more severe noise, it maintained robustness regarding ROUGE-L and METEOR metrics.
-Developing an even more robust framework is a goal for future research.
+5.5 Evaluation of Alignment with Human Expert Annotations
+In Table 4, we conducted an evaluation of model-generated ECG reports from ECG instruction-tuned versions of LLaMA-2 and LLaMA-3 against 500 ground-truth reports, meticulously annotated by human medical experts.
+These test annotated data were randomly sampled from the PTB-XL dataset, with all selected reports carefully reviewed and validated by human experts.
+Each model-generated report was compared with these expert-annotated reports using gpt-4o [^3] https://platform.openai.com/docs/models/gpt-4o, which assessed quality across four dimensions: Medical Terminology Accuracy, Logical Consistency, Completeness, and Diagnostic Accuracy, on a scale of 1 to 5.
+To evaluate these reports, we employed the following prompt template, which guided GPT-4o’s scoring process across the defined dimensions, as shown in Table 6 in Appendix.
+This prompt template ensures that GPT-4o evaluates the reports in a structured and consistent manner, highlighting both strengths and weaknesses of the model-generated reports in comparison to human expert annotations.
+The results indicate that the LLaMA-3 model, with an average Diagnostic Accuracy score of 3.85, closely matches the quality of the human expert annotations, whereas the LLaMA-2 model scored 3.60.
+This evaluation underscores the effectiveness of using human expert annotations from the PTB-XL (Wagner et al., 2020) dataset as a rigorous benchmark for assessing the models’ ability to generate clinically reliable ECG reports.
+6 Conclusion
+In this paper, we introduced MEIT, a framework for generating instruction-following data to train a multimodal LLM that can produce ECG reports based on human instructions.
+We also proposed an effective method for aligning ECG and report representations across various open-source LLMs, demonstrating strong performance on both the MIMIC-IV-ECG and PTB-XL datasets across multiple tasks.
+Additionally, we established a comprehensive benchmark for ECG instruction-following in report generation, providing a standardized evaluation for future research.
+Although this work primarily focuses on ECG signals, it serves as a foundational step in applying instruction-tuning to biomedical signals.
+For future research, we aim to extend our framework and benchmark to other medical domains, such as EEG, with the hope of driving further progress in developing more capable medical-signal LLMs.
+7 Limitations
+Our MEIT framework first attempts to address automatic ECG report generation through ECG instruction tuning, establishing the first comprehensive benchmark for this process with mainstream LLMs.
+In this work, we mainly focus on generating reports from multimodal ECG instructions using LLMs.
+However, the generated results are not fully explainable or controllable, even though the generation procedure is transparent and trackable.
+This is because the underlying theory of LLMs remains largely unexplored, necessitating further investigation to ensure the quality and safety of the generated content.
+In the future, we aim to enable LLMs to utilize external, expert-verified knowledge databases, such as clinical protocols and medical textbooks, to enhance the explainability of the generated ECG reports.
+8 Acknowledgment
+We thank the reviewers and ACs for their helpful comments.
+Appendix A Appendix.
+A.1 Further Analysis of MEIT
+Instruction Tuning Visualization.
+Figure 4 compares the convergence curves of the instruction tuning loss and the METEOR score between GPT-Neo (2.7B), BLOOM (7B), OPT (6.7B), and LLaMA-2 (7B) on the MIMIC-IV-ECG train and validation datasets.
+We observe that larger models with more parameters can converge to a more minor loss and achieve higher performance on the METEOR score.
+Notably, an increase in model size correlates with higher performance and lower loss, suggesting that larger models have the potential for better performance.
+Figure 4: Visualizations of instruction tuning loss and METEOR score.
 Table 5: Performance comparison of the proposed concatenated-fusion method and other mainstream fusion variants.
@@
 MEIT Concatenated-fusion 0.543 0.761 0.724 5.26
-5.3 Analysis
-Instruction Tuning Visualization.
-Figure 5 compares the convergence curves of the instruction tuning loss and the METEOR score between GPT-Neo (2.7B), BLOOM (7B), OPT (6.7B), and LLaMA-2 (7B) on the MIMIC-IV-ECG train and validation datasets.
-We observe that larger models with more parameters can converge to a more minor loss and achieve higher performance on the METEOR score.
-Notably, an increase in model size correlates with higher performance and lower loss, suggesting that larger models have the potential for better performance.
 Analysis of ECG Modality Alignment.
-To study the effectiveness of our proposed concatenated-fusion method for ECG modality alignment, we compare it with other fusion approaches such as direct input in LLaVA [33] and additional trainable cross-attention layer in Flamingo [12].
+To study the effectiveness of our proposed concatenated-fusion method for ECG modality alignment, we compare it with other fusion approaches such as direct input in LLaVA (Liu et al., 2023c) and additional trainable cross-attention layer in Flamingo (Alayrac et al., 2022).
 For straightforward input, we follow the design of LLaVA by directly concatenating the ECG encoder’s output embeddings with the sentence’s embeddings before inputting them into the LLM backbones.
@@
 Consequently, the concatenated fusion is more effective for the LLM backbone’s alignment with fine-grained ECG patterns without necessitating additional trainable parameters.
-Figure 5: Instruction tuning loss and METEOR visualization.
 Figure 5: Model scaling performance on MIMIC-IV-ECG and PTB-XL.
@@
 Similarly, on the MIMIC-IV-ECG dataset, the 70B model’s F-1 score is only 0.01 higher than that of the 13B model.
-Therefore, we conjecture that enhancing both data scale and model size concurrently is necessary to achieve superior performance [34].
+Therefore, we conjecture that enhancing both data scale and model size concurrently is necessary to achieve superior performance (Wei et al., 2022).
 Figure 6: Ablation Study of ECG Instruction Tuning on MIMIC-IV-ECG Dataset.
+Figure 7: Examples of ECG reports generated by LLaMA-2 and Mistral-Instruct. We highlight the consistent information between the generated reports and the ground truths with blue color.
 Ablation Study on ECG Instruction Tuning.
@@
 The findings, illustrated in Figure 6, indicate a significant performance drop across all metrics without instruction tuning, particularly in Mistral.
-This underscores instruction tuning’s superiority in enhancing LLMs’ generalization to new tasks/data over direct fine-tuning [35].
+This underscores instruction tuning’s superiority in enhancing LLMs’ generalization to new tasks/data over direct fine-tuning (Ouyang et al., 2022).
+Table 6: Prompt template used for GPT-4o evaluation. This prompt guided the model’s evaluation of generated ECG reports.
+Prompt Template for GPT-4o Evaluation
+You are an expert in Electrocardiogram (ECG) text evaluation. Your task is to assess the quality of a generated ECG report by comparing it to a real, expert-annotated ECG report.
+Generated ECG Report: {Generated_Report}
+Real ECG Report: {Real_Report}
+Please evaluate the generated report based on the following criteria:
+1. Medical Terminology Accuracy: Does the generated report use correct and appropriate ECG signal terms?
+2. Logical Consistency: Is the information presented in a logical and medically sound order?
+3. Completeness: Does the report include all necessary details that would be present in a real ECG report, such as heart rhythm, rate, and any abnormalities?
+4. Diagnostic Accuracy: Are the diagnoses and interpretations in the generated report accurate and consistent with the expert-annotated report?
+Please provide a detailed analysis and score each criterion on a scale of 1 to 5 (1 = Poor, 5 = Expert-Level).
 Qualitative Results.
@@
 These generated reports demonstrate the efficacy of our method.
-6 Related Works
-Instruction Tuning. Instruction tuning [36, 37] boosts zero-shot learning in LLMs for new tasks using instructions. Notable models like InstructGPT [35], FLAN-PaLM [38],
-Figure 8: Examples of ECG reports generated by LLaMA-2 and Mistral-Instruct. We highlight the consistent information between the generated reports and the ground truths with blue color.
-and Alpaca [39] fine-tune with instruction data through various methods, including human feedback.
-Similarly, multimodal models such as LLaVA [33], MiniGPT-4 [40], and AnyMAL [41] benefit from multimodal instructions for enhanced learning.
-However, instruction tuning for medical signals like ECG remains underexplored.
-Our research develops a targeted instruction-tuning framework for ECG report generation, addressing this gap.
-Medical Report Generation.
-Medical report generation [42]involves three strategies: (1) Template Selection and Generation, highlighted by HRGR [43] and CMAS [44]; (2) Data Integration and Coherence, as seen in PPKED [45] and CA [46]; (3) Cross-Modal Alignment, with efforts like [47, 48].
-While recent research has advanced these areas, applying them to ECG data is challenging due to the distinct nature of ECG signals compared to medical images.
-LLM for ECG Understanding.
-Only a few research efforts have focused on utilizing LLMs for ECG-related tasks [49, 50, 51].
-In the studies by [49, 51], the authors manually convert ECG signals into text features before feeding them into LLMs, rather than using the original ECG data.
-This approach overlooks the modality patterns inherent in the signals.
-Moreover, these studies only employ LLMs for classifying diseases from ECG data and do not explore the generation of medical reports.
-In contrast, [50] attempts ECG report generation by using handcrafted ECG features as inputs instead of the original signals.
-However, they have not publicly made any code or models available, making direct comparisons challenging.
-7 Conclusion and Future Work
-In this work, we propose and demonstrate the efficacy of the Multimodal ECG Instruction Tuning (MEIT) framework.
-Particularly, we first introduce a novel method for creating instruction-following data for ECG reports, enabling the training of an ECG report generator, a multimodal LLM designed to generate ECG reports following human instructions.
-Furthermore, we propose an effective and efficient method for aligning ECG and report representations across multiple open-source LLMs, achieving notable performance on both the MIMIC-IV-ECG and PTB-XL datasets for diverse tasks.
-Importantly, this paper introduces the comprehensive benchmark for ECG instruction-following in ECG report generation.
-While this research primarily concentrates on ECG signals, it marks a preliminary step in biomedical signal instruction tuning.
-In future work, we plan to extend our framework to encompass additional medical signal domains, such as EEG signals, with the aspiration that our work spurs further advancements in developing more proficient medical-signal LLMs.
-Appendix A Appendix.
-A.1 Limitations and Future Work
-In this work, we propose an effective and efficient method for generating ECG reports.
-Since this is a generative task, the content produced may occasionally lead to hallucinations.
-This is particularly pertinent in medical contexts such as ECG report generation, where accuracy is of utmost importance.
-To mitigate this issue, future iterations of our work could incorporate retrieval-augmented methods [14], information optimized strategies [52, 53, 54].
-For instance, by integrating a retrieval mechanism that accesses a database of verified medical or biological knowledge [55] or previous ECG reports, the model could be guided towards generating more accurate and reliable outputs.
 A.2 Hyper-parameters of ECG Instruction Tuning
@@
 Mistral-Instruct 128
-In this study, we implement the Low-Rank Adaptation (LoRA) [16] technique for efficient fine-tuning, specifically applied to ECG instruction tuning.
-As detailed in Table A.2 provided, we utilize mixed precision at bf16 for enhanced computational efficiency.
+In this study, we implement the Low-Rank Adaptation (LoRA) (Hu et al., 2021) technique for efficient fine-tuning, specifically applied to ECG instruction tuning.
+As detailed in Table 8 provided, we utilize mixed precision at bf16 for enhanced computational efficiency.
 Our models undergo instruction tuning over 5 epochs, with LoRA parameters set at an alpha of 64 and a rank of 128, accompanied by a dropout rate of 0.1.
@@
 We set the weight decay to 0.0.
-Moreover, as shown in Table A.2, we detail the ECG embedding dimensions for various language models, highlighting their approach to ECG data encoding.
+Moreover, as shown in Table 8, we detail the ECG embedding dimensions for various language models, highlighting their approach to ECG data encoding.
 GPT2-Medium and GPT2-Large feature ECG dimensions 64, while GPT-Neo, BLOOM, OPT, LLaMA-1, Mistral, LLaMA-2, and Mistral-Instruct use a dimension of 128.
@@
 A.3 More Details of ECG Encoder
-Projection Layer For the design of the projection layer within the ECG encoder, we adopt a non-linear approach similar to CLIP [56] and Med-UniC [57].
-Specifically, in our experiments, we employ two consecutive linear layers, each followed by BatchNorm1d [^4] https://pytorch.org/docs/stable/generated/torch.nn.BatchNorm1d.html.
+Projection Layer For the design of the projection layer within the ECG encoder, we adopt a non-linear approach similar to CLIP (Radford et al., 2021) and Med-UniC (Wan et al., 2024a).
+Specifically, in our experiments, we employ two consecutive linear layers, each followed by BatchNorm1d [^4] https://pytorch.org/docs/torch.nn.BatchNorm1d.html.
 Besides, ReLU serves as the activation function between the two linear layers.
@@
 The analysis reveals the ECG encoder’s trainable parameters are substantially fewer than those of the LoRA adapter in the LLM backbone during instruction tuning, and its parameter share of the overall framework is minimal for inference, underscoring its efficiency.
+Table 9: Comparisons of results with and without supervised manner. We take LLaMA-2-Instruct as the LLM backbone here. heavy teal color indicates the highest results.
+Methods Size MIMC-IV-ECG PTB-XL
+BLEU-4 METEOR ROUGE-L CIDEr-D MTA MTA LC DA
+Vision Mamba 86M 0.548 0.737 0.715 5.58 3.78 3.88 3.61 3.50
+Vision Transformer 98M 0.592 0.815 0.772 5.67 4.33 4.15 4.12 3.78
+Vision Transformer (SSL) 98M 0.581 0.822 0.766 5.75 4.42 4.28 3.85 3.85
+1-D Temporal Conv (Ours) 20.4M 0.610 0.799 0.773 5.78 4.52 4.38 4.01 3.98
+Ablation Study of ECG Encoder we conducted additional experiments comparing our default 1-D Temporal Convolution ECG encoder with alternative architectures, including: 1.
+S4-based Model: Vim-B (Vision Mamba, 98M parameters) (Zhu et al., 2024a).
+2.
+Transformer-based Model: ViT-B/16 (Vision Transformer, 86M parameters) (Dosovitskiy et al., 2020), adapted for 1-D token patching to align with the temporal nature of ECG signals.
+3.
+SSL-Transformer Model: ViT-B/75 initialized with self-supervised learning (SSL) weights specific to ECG signals (Na et al., 2024).
+We evaluated these models on two tasks: Quality of Generated Reports using the MIMIC-IV-ECG dataset, and Evaluation of Alignment with Human Expert Annotations using the PTB-XL dataset.
+For fair comparison, we used Meta-Llama-3-8B-Instruct as the LLM backbone due to its consistent strong performance.
+The results, summarized in the table below, show that our 1-D Temporal Convolution ECG encoder, despite having significantly fewer parameters, performs comparably or better across most metrics compared to ViT and ViT-SSL, and comprehensively outperforms the S4-based Vim.
+Notably, the ViT-SSL encoder demonstrates the benefit of self-supervised pretraining for initial ECG representation learning.
+However, our default ECG encoder effectively captures the 12-channel ECG temporal patterns while remaining lightweight, making it well-suited for our efficient instruction tuning framework.
+These findings validate the effectiveness of our 1-D Temporal Conv encoder and also provide valuable insights for future work, including designing more complex ViT-based architectures optimized for ECG time-series data.
 Table 10: Parameter Comparison of ECG encoder and LLM backbone. We use LLaMA-1 7B as an example.
@@
 A.4 Further Analysis of Generated Prompts
-In the ECG instruction data curation, we manually created 32 prompt examples, as illustrated in Section 3.1.
+Prompts Number Analysis In the ECG instruction data curation, we manually created 32 prompt examples, as illustrated in Section 3.3.
 To increase the diversity of our samples, we employed GPT-4 to rephrase these manually designed prompts, generating a larger pool of prompt examples.
@@
 512 0.538 0.754 0.732 5.03
-A.5 Visualization of Generated ECG Report Samples
-As illustrated in Figures 9,  10, and  11 we have visualized the report samples generated by LLaMA-1, LLaMA-2, and Mistral-Instruct.
+Ablation Study on GPT-4 Prompt Rephrasing We also conducted an ablation study to compare the performance with and without GPT-4 rephrasing prompts, using a fixed prompt for the latter.
+The results in the following Table 12 indicate that using diverse prompts rephrased by GPT-4 leads to better performance, highlighting the superiority of instruction tuning in enhancing LLMs’ generalization to new tasks and data over direct fine-tuning.
+Table 12: Performance comparison of with and without GPT-4 prompt rephrasing. We take Mistral-Instruct as the LLM backbone here. heavy teal color indicates the highest results.
+Prompt Nums BLEU-4 METEOR ROUGE-L CIDEr-D
+w.o. Rephrasing 0.564 0.745 0.738 5.50
+w. Rephrasing (Ours) 0.576 0.768 0.751 5.62
+A.5 Comparison with Encoder-Decoder Models
+In this section, we conducted additional comparative experiments using two open-source traditional encoder-decoder architectures: BART-Large (406M parameters) (Lewis, 2019) and T5-Large (780M parameters) (Raffel et al., 2020), as shown in Table 13.
+In adapting our framework for ECG instruction tuning, we employ the language encoder to process the input instruction, an ECG encoder to handle the input ECG signals, and the language decoder to generate the ECG report based on the output from both language end ECG encoder.
+Our findings indicate that the performance of encoder-decoder models is comparable to the small pre-trained language models (GPT2-Medium and GPT-Large) presented in Table 1 and Table 2 of our paper.
+Moreover, LLM-based backbones (such as LLaMA1-2) consistently achieve a significant margin of improvement over the encoder-decoder architectures across all metrics.
+Table 13: Comparison with encoder-decoder-based models on MIMIC-IV-ECG.
+For model size, ’M’ denotes the million level, and ’B’ denotes the billion level.
+The light teal color indicates the second highest results, and heavy teal color indicates the highest results.
+Models Size BLEU-1 BLEU-2 BLEU-3 BLEU-4 METEOR ROUGE-L ROUGE-1 ROUGE-2 CIDEr-D
+BART-Large 406M 0.525 0.498 0.466 0.388 0.455 0.472 0.5124 0.451 3.15
+T5-Large 780M 0.595 0.542 0.465 0.422 0.498 0.456 0.522 0.438 4.08
+LLaMA-1 7B 0.685 0.648 0.615 0.543 0.761 0.724 0.742 0.642 5.26
+LLaMA-2-Instruct 7B 0.706 0.662 0.622 0.581 0.775 0.745 0.768 0.664 5.55
+A.6 Analysis of Combining MEIT with a Supervised Manner
+In this section, we conduct a new experiment where we trained a CNN (ECG encoder) in a supervised manner on the PTB-XL training set, utilizing all available annotations (approximately 70 patterns), as shown in Table 14.
+We then transferred the CNN for ECG instruction fine-tuning on both the MIMIC-IV-ECG and PTB-XL datasets.
+Our findings indicate that performance increased on the PTB-XL dataset in most metrics, likely due to the model’s prior learning of specific annotated patterns.
+However, performance fluctuated on the MIMIC-IV-ECG dataset, which contains more data and exhibits greater diversity.
+This suggests that the supervised approach may enhance performance on in-domain data, but it limits generalizability to data from unseen domains.
+Table 14: Comparisons of results with and without supervised manner. We take LLaMA-2-Instruct as the LLM backbone here. heavy teal color indicates the highest results.
+Methods PTB-XL
+BLEU-4 METEOR ROUGE-L CIDEr-D
+MEIT 0.439 0.675 0.594 4.05
+MEIT + Supervised manner 0.445 0.664 0.612 4.12
+MIMIC-IV-ECG
+BLEU-4 METEOR ROUGE-L CIDEr-D
+MEIT 0.581 0.775 0.745 5.55
+MEIT + Supervised manner 0.578 0.778 0.739 5.47
+A.7 Computational Cost Analysis of MEIT
+The time cost experiment, detailed in the Table 15, was conducted on the MIMIC-IV-ECG dataset.
+We found that larger models have longer training and inference times.
+Thus, we are considering techniques like quantization and other compression methods to improve model efficiency in future work.
+Table 15: Computational time Analysis of MEIT with various parameters and backbones.
+Model Size Training time Testing time
+4 A100 and 3 Epochs 1 A100 and 128 Generated Samples
+GPT-2 Large 774M 3.25h 3.125min
+LLaMA-2-Instruct 7B 13.5h 9 min
+LLaMA-2-Instruct (+) 13B 27h 14.125 min
+A.8 Examples of ECG Instruction Tuning Data and the Corresponding Generated ECG Reports
+As illustrated in Figures 8,  9, and  10 we have visualized the report samples generated by LLaMA-1, LLaMA-2, and Mistral-Instruct.
 The samples are presented in blue font to highlight the key information that aligns with the ground truth.
@@
 Figure 10: Reports generated by Mistral-Instruct following ECG instruction Tuning.
+A.9 Future Works
+In this work, we propose an effective and efficient method for generating ECG reports.
+Since this is a generative task, the content produced may occasionally lead to hallucinations Zhu et al.
+(2025).
+This is particularly pertinent in medical contexts such as ECG report generation, where accuracy is of utmost importance.
+To mitigate this issue, future iterations of our work could incorporate retrieval-augmented methods Wan et al.
+(2022); Li et al.
+(2024), information optimized strategies Xiong et al.
+(2022); Liang et al.
+(2020a, b), multi-modal fusion and optimization strategies Wan et al.
+(2024c); Xiong et al.
+(2024a); Liu et al.
+(2024c); Gong et al.
+(2024); Wan et al.
+(2023a); Chen et al.
+(2025); Luo et al.
+(2024); Shen et al.
+(2025b); Wan et al.
+(2025b); Zhang et al.
+(2025b); Liu et al.
+(2024b); Huang et al.
+(2024); Zhu et al.
+(2024b); Liu et al.
+(2025a), Reasoning enhanced methods Wan et al.
+(2025a); Liu et al.
+(2025b); Shen et al.
+(2025a), efficient ML Wang et al.
+(2024b); Tao et al.
+(2024); Wan et al.
+(2024b); Liu (2024); Xin et al.
+(2024); Shen et al.
+(2024); Wang et al.
+(2025); Xiong et al.
+(2024b, 2025).
+For instance, by integrating a retrieval mechanism that accesses a database of verified medical or biological knowledge Zheng et al.
+(2024) or previous ECG reports, the model could be guided towards generating more accurate and reliable outputs.